### v2

# Fibred sets within a predicative and constructive effective topos

Abstract
We describe the fibrational structure of sets within the predicative variant $\mathbf{pEff}$ of Hyland’s Effective Topos ${\mathbf{Eff}}$ previously introduced in Feferman’s predicative theory of non-iterative fixpoints $\widehat{ID_{1}}$. Our structural analysis can be carried out in constructive and predicative variants of ${\mathbf{Eff}}$ within extensions of Aczel’s Constructive Zermelo-Fraenkel Set Theory. All this shows that the full subcategory of discrete objects of Hyland’s Effective topos ${\mathbf{Eff}}$ contains already a fibred predicative topos validating the formal Church’s thesis, even when both are formalized in a constructive metatheory.
1 Introduction
This paper builds on the work of [MM21] by further contributing to a predicative and constructive development of elementary topos theory, drawing inspiration from [Hof04, MP02, BM11, BM08].
More in detail, we investigate the fibred structure of sets of the predicative version $\mathbf{pEff}$ of Hyland’s “Effective Topos” in [MM21], said "predicative" since $\mathbf{pEff}$ is formalized in Feferman’s weak theory of inductive definitions $\widehat{ID_{1}}$ [Fef82].
Here, in addition, we also show that our fibred $\mathbf{pEff}$ can be formalized in a constructive and predicative meta-theory, such as Aczel’s $\mathbf{CZF}+{\bf REA_{\bigcup}}+\bf RDC$, so that inductive and coinductive predicates can be defined in its internal logic.
The ultimate goal of this investigation is to show that $\mathbf{pEff}$ has a fibred categorical structure sufficient to model the predicative and constructive two-level Minimalist Foundation, for short MF, conceived in [MS05] and finalized in [Mai09], and its extensions with (co)-inductive predicates in [MMR21, MMR22, MS23, Sab24]. In this way, $\mathbf{pEff}$ would provide a categorical predicative and constructive universe where to exhibit the constructive contents of proofs performed in MF and extensions in terms of programs. All this would contribute to broadening further the already wide range of applications of Hyland’s “Effective Topos” ${\mathbf{Eff}}$ in [Hyl82], whose introduction opened different research lines connecting computability to topos theory with new applications to logic, mathematics, and computer science.
A main reason for such a broad range of applications is that ${\mathbf{Eff}}$ provides a realizability interpretation of higher-order logic extending Kleene realizability for intuitionistic arithmetic and, hence, validating formal Church’s thesis, (see [vO94]). Thanks to the results in this paper, our $\mathbf{pEff}$ would become a predicative realizability model for MF  extended with inductive and coinductive predicates and formal Church’s thesis.
A key aspect of ${\mathbf{Eff}}$ that will help to reach our goal is that the structure of ${\mathbf{Eff}}$ roughly resembles the two-level structure of MF in [Mai09], which is one of the peculiarities of MF among the variety of foundations for mathematics (and it is crucial to show its minimality by comparing each foundation with the most appropriate level of MF).
More in detail, the two-level structure of MF in [Mai09] consists of an intensional level that can be used as a base for a proof-assistant and extraction of computational contents from proofs, an extensional level where to develop mathematics in a language close to the usual practice, and an interpretation of the extensional level in the intensional one using a quotient model. Such a quotient model was analyzed categorically in [MR13b, MR13a] as an instance the elementary quotient completion of a Lawvere’s elementary doctrine introduced in loc.cit as a generalization of the well-known notion of exact completion on a lex category in [Car95, CM82, CV98].
Analogously, the structure of ${\mathbf{Eff}}$ is an instance of a quotient completion called tripos-to-topos construction introduced in [HJP80] (see [MR15, MR16] for a precise comparison with the notion of elementary quotient completion). But, while ${\mathbf{Eff}}$ enjoys an impredicative internal language that can be modeled by just using suitable Lawvere doctrines, called triposes, both levels of MF are formulated as constructive and predicative dependent type theories which need to be modeled by using fibrations or their variants as in [Jac99, Awo18, Dyb96, Mai05].
Even the predicative version $\mathbf{pEff}$ of ${\mathbf{Eff}}$ in [MM21] was built as the elementary quotient completion of a categorical rendering of the interpretation of the intensional level of MF within Feferman’s $\widehat{ID_{1}}$ described in [IMMS18]. What remains is to model the extensional level emTT of MF in $\mathbf{pEff}$. Now, as hyperdoctrines are just enough to model the first order logical part of MF, to model also the set-theoretic part of MF, and in particular of emTT, we need to enucleate the fibred structure of $\mathbf{pEff}$ by analyzing its codomain fibration.
More in detail, since both levels of MF  distinguish among collections, sets, propositions (quantifying on arbitrary collections), and small propositions (quantifying only on sets), to model all of them we consider an equivalent presentation of the codomain fibration over $\mathbf{pEff}$ whose fibers are families of dependent collections, which are in turn equivalent to suitable internal groupoid actions in the sense of [MM12, JT94] within $\mathbf{pEff}$. Then, we identify a subfibration of families of sets of the codomain fibration of $\mathbf{pEff}$, intended to model respectively sets and collections of MF, whose corresponding sub-fibrations of subobjects will be used to model small propositions and propositions of MF.
${{{\mathbf{pEff}^{\to}_{set}}}}$${{{\mathbf{pEff}}^{\to}}}$${{\mathbf{pEff}}}$$\scriptstyle{\mathsf{cod}}$$\scriptstyle{\mathsf{cod}}$
Such a fibred structure can be shown to exists also when $\mathbf{pEff}$ is formalized in $\mathbf{CZF}$ + $\mathbf{REA}_{\bigcup}+\mathbf{RDC}$ by defining it as the elementary quotient completion of a categorical rendering of the realizability interpretation of the intensional level of MF enriched with inductive and coinductive topological predicates in [MMR21, MMR22]. So, $\mathbf{pEff}$ would gain the expressive power to model both levels of MF extended with inductive and coinductive predicates thanks to the results in [MS23, Sab24].
Furthermore, as already observed in [MM21], our constructions of $\mathbf{pEff}$ can be mapped into ${\mathbf{Eff}}$, and actually in its subcategory of the exact on lex completion of recursive objects, called discrete objects in [HRR90, vO15], thanks to the fact that ${\mathbf{Eff}}$ can be viewed as the exact on lex completion of the category of partitioned assemblies (see [RR90, MT22, MT21]).
It is worth recalling that a constructive predicative study of ${\mathbf{Eff}}$ was developed in the context of algebraic set theory by B. van den Berg and I. Moerdijk (see [BM11]) by taking Aczel’s $\mathbf{CZF}$ in [AR01] as the set theory to be realized in such toposes. The full subcategory of $\mathbf{pEff}$ consisting of objects equipped small equivalence relations can be embedded into van den Berg and Moerdijk’s predicative version of ${\mathbf{Eff}}$, when formalized in $\mathbf{CZF}$ + $\mathbf{REA}_{\bigcup}+\mathbf{RDC}$. Instead, the whole constructive and predicative $\mathbf{pEff}$ can only be embedded into the category of discrete objects [HRR90, vO15] in the same metatheory.
From these comparisons we conclude the full subcategory of discrete objects of Hyland’s Effective topos ${\mathbf{Eff}}$ contains already a fibred predicative topos validating the formal Church’s thesis, even when both are formalized in a constructive metatheory.
As a future goal, recalling that ${\mathbf{Eff}}$ was built as a first example of a non-Grothendieck topos sharing with localic toposes the structure of tripos-to-topos construction, we hope to exploit a general notion of fibred predicative elementary topos and a predicative tripos-to-topos construction to include predicative presentations of realizability and localic toposes. Thanks to the results in this paper, our fibred $\mathbf{pEff}$ constitutes a key example for pursuing this project.
2 Our predicative meta-theory.
2.1 A classical predicative metatheory l̀a Feferman: $\widehat{ID_{1}}$.
Feferman’s theory of non-iterative fixpoints $\widehat{ID_{1}}$ [Fef82] is a first-order classical theory with equality containing 00, $succ$, the functional symbols for the (definitions of) primitive recursive functions and a unary predicate symbol $P_{\varphi}$ for every admissible second-order formula $\varphi(x,X)$ (see [Fef82, MM16, MM21] for more details).
The axioms of $\widehat{ID_{1}}$ include those of Peano arithmetic (even with those defining equations for primitive recursive functions) plus the induction principle for every formula of $\widehat{ID_{1}}$ and a fixpoint schema: $P_{\varphi}(x)\leftrightarrow\varphi[P_{\varphi}/X]$ for every admissible second-order formula $\varphi$, where $\varphi[P_{\varphi}/X]$ is the result of substituting in $\varphi$ all the subformulas of the form $t\,\epsilon\,X$ with $P_{\varphi}(t)$.
We follow the notation for recursive functions in $\widehat{ID_{1}}$ in [MM21].
2.2 A constructive and predicative meta-theory: extensions of CZF
P. Aczel introduced $\mathbf{CZF}$ as a constructive theory of sets (see [AR01] for its axioms). In this paper we are interested in two extensions of it used in [MMR21] and [MMR22] to prove the consistency of the Minimalist Foundation with inductive and coinductive topological generation, respectively. The first extension adds to $\mathbf{CZF}$ the so-called regular extension axiom $\mathbf{REA}$ which states that every set is a subset of a regular set. The second extension adds to $\mathbf{CZF}$ the axiom $\bigcup-\mathbf{REA}$, stating that every set is contained in a $\bigcup$-regular set, and the axiom of relativized dependent choice ($\mathbf{RDC}$). In [MMR21] and [MMR22] one can see how these axioms play a crucial role in interpreting universes and inductive and coinductive topological generation. Further details on these theories can be found in [AR01].
In what follows we will write $\mathcal{T}$ as a metavariable for any of the theories $\widehat{ID_{1}}$, $\mathbf{CZF}+\mathbf{REA}$ or $\mathbf{CZF}+\mathbf{RDC}+\bigcup-\mathbf{REA}$.
3 The starting structure of realized collections, sets, and propositions
From [MM21], we recall here the definition of the category of realized collections called $\bm{\mathcal{C}}_{r}$ together with four indexed categories defined over it: that of realized collections and its sub-indexed category of realized sets
${{\bm{\mathcal{C}}_{r}^{\operatorname{op}}}}$${{\mathbf{Cat}}}$${{\bm{\mathcal{C}}_{r}^{\operatorname{op}}}}$$\scriptstyle{\mathbf{Set}^{r}}$$\scriptstyle{id}$$\scriptstyle{\mathbf{Col}^{r}}$
and the doctrine of realized proposition and its sub-doctrine of small propositions
${{\bm{\mathcal{C}}_{r}^{\operatorname{op}}}}$${{\mathbf{Pos}}}$${{\bm{\mathcal{C}}_{r}^{\operatorname{op}}}}$$\scriptstyle{\overline{{\mathbf{Prop}^{r}_{\mathbf{s}}}}}$$\scriptstyle{id}$$\scriptstyle{\overline{{\mathbf{Prop}^{r}}}}$
Then, we consider the predicative effective topos $\mathbf{pEff}$ in [MM21] defined as the base $\mathcal{Q}_{\overline{\mathbf{Prop}^{r}}}$ of the elementary quotient completion of $\overline{{\mathbf{Prop}^{r}}}$.
$\mathbf{pEff}:\equiv(\mathcal{Q}_{\overline{\mathbf{Prop}^{r}}})$
The key point of this paper is to show how to lift the indexed category $\mathbf{Set}^{r}$, and consequently also $\overline{{\mathbf{Prop}^{r}_{\mathbf{s}}}}$, on the predicative effective topos $\mathbf{pEff}$ to the aim of investigating their categorical properties. We will do this by switching from indexed categories to fibrations.
3.1 The category $\bm{\mathcal{C}}_{r}$ of realized collections of $\mathcal{T}$.
A realized collection (or simply a class) $\mathsf{A}$ of $\mathcal{T}$ is a formal expression $\{x|\,\varphi_{\mathsf{A}}\}$ representing a subclass of natural numbers in $\mathcal{T}$ [^1] In the set theoretical case, this means that $\varphi_{\mathsf{A}}(x)\vdash_{\mathcal{T}}Nat(x)$ where $Nat(x)$ is the formula expressing the fact that $x$ is a natural number in $\mathcal{T}$. where $\varphi_{\mathsf{A}}$ is a formula of $\mathcal{T}$ with at most $x$ as free variable. We write $x\,\varepsilon\,\mathsf{A}$ as an abbreviation for $\varphi_{\mathsf{A}}$. Classes with provably equivalent membership relations in $\mathcal{T}$ are identified. An operation between classes of $\mathcal{T}$ from $\mathsf{A}$ to $\mathsf{B}$ is an equivalence class $[\mathbf{n}]_{\approx_{\mathsf{A},\mathsf{B}}}$ of numerals with $x\,\varepsilon\,\mathsf{A}\vdash\{\mathbf{n}\}(x)\,\varepsilon\,\mathsf{B}$ in $\mathcal{T}$. For such $\mathbf{n}$ and $\mathbf{m}$ we define $\mathbf{n}\approx_{\mathsf{A},\mathsf{B}}\mathbf{m}$ as $x\,\varepsilon\,\mathsf{A}\vdash\{\mathbf{n}\}(x)=\{\mathbf{m}\}(x)$ in $\mathcal{T}$.
If $[\mathbf{n}]_{\approx_{\mathsf{A},\mathsf{B}}}:\mathsf{A}\rightarrow\mathsf{B}$ and $[\mathbf{m}]_{\approx_{\mathsf{B},\mathsf{C}}}:\mathsf{B}\rightarrow\mathsf{C}$ are operations between classes of $\mathcal{T}$, then their composition is defined by $[\mathbf{m}]_{\approx_{\mathsf{B},\mathsf{C}}}\circ[\mathbf{n}]_{\approx_{\mathsf{A},\mathsf{B}}}:=[\Lambda x.\{\mathbf{m}\}(\{\mathbf{n}\}(x))]_{\approx_{\mathsf{A},\mathsf{C}}}$.
The identity operation for the class $\mathsf{A}$ of $\mathcal{T}$ is defined as $\mathsf{id}_{\mathsf{A}}:=[\Lambda x.x]_{\approx_{\mathsf{A},\mathsf{A}}}$.
Definition 3.1.
We denote with $\bm{\mathcal{C}}_{r}$ the category whose objects are realized collections of $\mathcal{T}$ and arrows are operations in $\mathcal{T}$ between them with their composition and identity operations.
The following result appears in [MM21, Theorem 4.3].
Theorem 3.2.
$\bm{\mathcal{C}}_{r}$ is a finitely complete category with disjoint stable finite coproducts, parameterized list objects, and weak exponentials. Moreover, finite coproducts in $\bm{\mathcal{C}}_{r}$ are disjoint and stable. ∎
3.2 Families of realized collections of $\mathcal{T}$.
Here, we provide a functorial account of the slice pseudofunctor over the the category of collection $\bm{\mathcal{C}}_{r}$ by defining an indexed category
$\mathbf{Col}^{r}:\bm{\mathcal{C}}_{r}^{op}\rightarrow\mathbf{Cat}.$ (1)
as follows.
Suppose $\mathsf{A}$ is an object of $\bm{\mathcal{C}}_{r}$. A family of realized collections $\mathsf{C}$ on $\mathsf{A}$ is a formal expression $\{x^{\prime}|\,\varphi_{\mathsf{C}}\}$ representing a subclass of natural numbers depending on natural numbers in $\mathcal{T}$ [^2] In the set theoretical case, this means that $\varphi_{C}\vdash_{\mathcal{T}}Nat(x)\wedge Nat(x^{\prime})$. where $\varphi_{\mathsf{C}}$ is a formula of $\mathcal{T}$ with at most $x$ and $x^{\prime}$ as free variables (we write $x^{\prime}\,\varepsilon\,\mathsf{C}$ as an abbreviation for $\varphi_{\mathsf{C}}$) for which $x^{\prime}\,\varepsilon\,\mathsf{C}\,\vdash\,x\,\varepsilon\,\mathsf{A}$ in $\mathcal{T}$. Families of realized collections on $\mathsf{A}$ with provably (in $\mathcal{T}$) equivalent membership relations are identified.
An operation from a family of realized collections $\mathsf{C}$ on $\mathsf{A}$ to another $\mathsf{D}$ is given by an equivalence class $[\mathbf{n}]_{\approx_{\mathsf{C},\mathsf{D}}}$ of numerals such that in $\mathcal{T}$ we have $x^{\prime}\,\varepsilon\,\mathsf{C}\,\vdash\,\{\mathbf{n}\}(x,x^{\prime})\,\varepsilon\,\mathsf{D}$ with respect to the equivalence relation defined as follows: $\mathbf{n}\approx_{\mathsf{C},\mathsf{D}}\mathbf{m}$ if and only if $x^{\prime}\,\varepsilon\,\mathsf{C}\,\vdash_{\mathcal{T}}\,\{\mathbf{n}\}(x,x^{\prime})=\{\mathbf{m}\}(x,x^{\prime})$.
If $\mathsf{f}=[\mathbf{n}]_{\approx_{\mathsf{C},\mathsf{D}}}:\mathsf{C}\rightarrow\mathsf{D}$ and $\mathsf{g}=[\mathbf{m}]_{\approx_{\mathsf{D},\mathsf{E}}}:\mathsf{D}\rightarrow\mathsf{E}$ are operations between families of realized collection on $\mathsf{A}$, then their composition is defined as
$\mathsf{g}\circ\mathsf{f}:=[\Lambda x.\Lambda x^{\prime}.\{\mathbf{m}\}(x,\{\mathbf{n}\}(x,x^{\prime}))]_{\approx_{\mathsf{C},\mathsf{E}}}:\mathsf{C}\rightarrow\mathsf{E}.$
If $\mathsf{C}$ is a family of realized collections on $\mathsf{A}$, then its identity operation is defined by $\mathsf{id}_{\mathsf{C}}:=[\Lambda x.\Lambda x^{\prime}.x^{\prime}]_{\approx_{\mathsf{C},\mathsf{C}}}:\mathsf{C}\rightarrow\mathsf{C}$.
As discussed in [MM21, Theorem 4.6], for every object $\mathsf{A}$ of $\bm{\mathcal{C}}_{r}$, families of realized collections on $\mathsf{A}$ and operations between them together with their composition and identity operations define a category denoted with $\mathbf{Col}^{r}(\mathsf{A})$.
If $\mathsf{f}:=[\mathbf{n}]_{\approx}:\mathsf{A}\rightarrow\mathsf{B}$, then the following assignments give rise to a functor $\mathbf{Col}^{r}_{\mathsf{f}}$ from $\mathbf{Col}^{r}(\mathsf{B})$ to $\mathbf{Col}^{r}(\mathsf{A})$:
  1. 1.
    for every object $\mathsf{C}$ of $\mathbf{Col}^{r}(\mathsf{B})$, $\mathbf{Col}^{r}_{\mathsf{f}}(\mathsf{C}):=\{x^{\prime}|\,x\,\varepsilon\,\mathsf{A}\ \wedge\ x^{\prime}\,\varepsilon\,\mathsf{C}[\{\mathbf{n}\}(x))/x]\ \}$;
  2. 2.
    for every arrow $\mathsf{g}:=[\mathbf{k}]_{\approx}:\mathsf{C}\rightarrow\mathsf{D}$ of $\mathbf{Col}^{r}(\mathsf{B})$,
    $\mathbf{Col}^{r}_{\mathsf{f}}(\mathsf{g}):=[\Lambda x.\Lambda x^{\prime}.\{\mathbf{k}\}(\{\mathbf{n}\}(x),x^{\prime})]_{\approx}:{\mathbf{Col}^{r}_{\mathsf{f}}(\mathsf{C})\rightarrow\mathbf{Col}^{r}_{\mathsf{f}}(\mathsf{D})}.$
Moreover, the assignments $\mathsf{A}\mapsto\mathbf{Col}^{r}(\mathsf{A})$ and $\mathsf{f}\mapsto\mathbf{Col}^{r}_{\mathsf{f}}$ define an indexed category
$\mathbf{Col}^{r}:\bm{\mathcal{C}}_{r}^{op}\rightarrow\mathbf{Cat}.$ (2)
The functors $\mathbf{Col}^{r}_{\mathsf{f}}$ are also called substitution functors and, as proven in [MM21, Theorem 4.8], these functors have left adjoints
$\Sigma_{\mathsf{f}}:\mathbf{Col}^{r}(\mathsf{A})\rightarrow\mathbf{Col}^{r}(\mathsf{B})$ (3)
which sends an object $\mathsf{C}\in\mathbf{Col}^{r}(\mathsf{A})$ to
$\Sigma_{\mathsf{f}}(\mathsf{C}):=\{x^{\prime}|x=\{\mathbf{n}\}(p_{1}(x^{\prime}))\wedge\ p_{2}(x^{\prime})\ \varepsilon\ \mathsf{C}[p_{1}(x^{\prime})/x]\}$
and an arrow $\mathsf{g}:=[\mathbf{m}]_{\approx}:\mathsf{C}\to\mathsf{D}$ to
$\Sigma_{\mathsf{f}}(\mathsf{g}):=[\Lambda x.\Lambda x^{\prime}.\{\mathbf{pair}\}(\{\mathbf{p}_{1}\}(x^{\prime}),\{\mathbf{m}\}(\{\mathbf{p}_{1}\}(x^{\prime}),\{\mathbf{p}_{2}\}(x^{\prime})))]_{\approx}$
Furthermore, in [MM21, Theorem 4.9] the authors provide also a weak versions of right adjoints
$\Pi_{\mathsf{f}}:\mathbf{Col}^{r}(\mathsf{A})\rightarrow\mathbf{Col}^{r}(\mathsf{B})$ (4)
Now we recall an equivalent presentation of families of realized collections, the following result appears as [MM21, Theorem 4.12].
Theorem 3.3.
For every $\mathsf{A}$ in $\bm{\mathcal{C}}_{r}$, $\bm{\mathcal{C}}_{r}/\mathsf{A}$ and $\mathbf{Col}^{r}(\mathsf{A})$ are equivalent. ∎
Therefore $\mathbf{Col}^{r}$ is a functorial account of the slice pseudofunctor on $\bm{\mathcal{C}}_{r}$ and hence it provides a split fibration
$\pi:\mathsf{Gr}(\mathbf{Col}^{r})\to\bm{\mathcal{C}}_{r}$
3.3 Families of realized sets
[…]
Definition 3.4.
Let $\mathsf{A}$ be an object of $\bm{\mathcal{C}}_{r}$. The category $\mathbf{Set}^{r}(\mathsf{A})$ is the full subcategory of $\mathbf{Col}^{r}(\mathsf{A})$
${{\mathbf{Set}^{r}(\mathsf{A})}}$${{\mathbf{Col}^{r}(\mathsf{A})}}$
whose objects, called families of realized sets on $\mathsf{A}$, are families of realized collections on $\mathsf{A}$ of the form
$\tau_{\mathsf{A}}(\mathbf{n}):=\{x^{\prime}|\,x^{\prime}\,\overline{\varepsilon}\,\{\mathbf{n}\}(x)\,\wedge\,x\,\varepsilon\,\mathsf{A}\}$
for a numeral $\mathbf{n}$ defining an operation $[\mathbf{n}]_{\approx}:\mathsf{A}\rightarrow\mathsf{U_{S}}$ in $\bm{\mathcal{C}}_{r}$.
Applying the Grothendieck construction we obtain the diagram
${{{\mathsf{Gr}(\mathbf{Set}^{r})}}}$${{{\mathsf{Gr}(\mathbf{Col}^{r})}}}$${{\bm{\mathcal{C}}_{r}}}$$\scriptstyle{\pi}$$\scriptstyle{\pi}$
3.4 Categorical properties of realized collections and sets
We recall the properties of the categories of families of realized collections $\mathbf{Col}^{r}(\mathsf{A})$. The following result appears as [MM21, Theorem 4.7].
Theorem 3.5.
For every object $\mathsf{A}$ of $\bm{\mathcal{C}}_{r}$, $\mathbf{Col}^{r}(\mathsf{A})$ is a finitely complete category with parameterized list objects, finite coproducts, and weak exponentials. Moreover for every arrow $\mathsf{f}:\mathsf{A}\rightarrow\mathsf{B}$ in $\bm{\mathcal{C}}_{r}$, the functor $\mathbf{Col}^{r}_{\mathsf{f}}$ preserves this structure. ∎
We recall the main ingredients of the proof of Theorem 3.3 which will be used in the rest of the paper. The proof of the above result relies on the definition of the following functors. The first one is
$\mathbf{J}_{\mathsf{A}}:\bm{\mathcal{C}}_{r}/\mathsf{A}\rightarrow\mathbf{Col}^{r}(\mathsf{A})$ (5)
defined by the assignments
$\begin{array}[]{rcl}[\mathbf{b}]_{\approx}:\mathsf{B}\rightarrow\mathsf{A}&\mapsto&\{x^{\prime}|\,x^{\prime}\,\varepsilon\,\mathsf{B}\wedge x=\{\mathbf{b}\}(x^{\prime})\}\\[3.0pt] [\mathbf{n}]_{\approx}&\ \mapsto&\ [\Lambda x.\Lambda x^{\prime}.\{\mathbf{n}\}(x^{\prime})]_{\approx}.\end{array}$
The second is the functor
$\mathbf{I}_{\mathsf{A}}:\mathbf{Col}^{r}(\mathsf{A})\rightarrow\bm{\mathcal{C}}_{r}/\mathsf{A}$ (6)
defined by the assignments
$\begin{array}[]{rcl}\mathsf{B}&\mapsto&\ \mathsf{p}_{1}^{\Sigma}:=[\mathbf{p}_{1}]_{\approx}:\Sigma(\mathsf{A},\mathsf{B})\rightarrow\mathsf{A}\\[3.0pt] [\mathbf{n}]_{\approx}&\mapsto&\ [\Lambda x.\{\mathbf{pair}\}(\{\mathbf{p}_{1}\}(x),\{\mathbf{n}\}(\{\mathbf{p}_{1}\}(x),\{\mathbf{p}_{2}\}(x)))]_{\approx}.\end{array}$
where $\Sigma(\mathsf{A},\mathsf{B})$ is the object defined as
$\Sigma(\mathsf{A},\mathsf{B}):=\{x|\,p_{1}(x)\,\varepsilon\,\mathsf{A}\wedge p_{2}(x)\,\varepsilon\,\mathsf{B}[p_{1}(x)/x])\}.$ (7)
The above Theorem implies the following result which appears as [MM21, Corollary 4.13].
Corollary 3.6.
$\bm{\mathcal{C}}_{r}$ is weakly locally cartesian closed.
Observe that, since every arrow in $\bm{\mathcal{C}}_{r}$ can be represented by many different numerals, the category $\bm{\mathcal{C}}_{r}$ is weakly locally cartesian closed but not cartesian closed.
Now we recall the properties of the category of families of realized sets.
Definition 3.7.
An arrow $\mathsf{f}:\mathsf{A}\rightarrow\mathsf{B}$ in $\bm{\mathcal{C}}_{r}$ is set-fibred if there exists an object $\mathsf{C}$ in $\mathbf{Set}^{r}(\mathsf{B})$ for which $\mathsf{f}$ is isomorphic in $\bm{\mathcal{C}}_{r}/\mathsf{B}$ to $\mathbf{I}_{\mathsf{B}}(\mathsf{C})$.
Lemma 3.8.
If $\mathsf{f}:\mathsf{B}\rightarrow\mathsf{C}$ is an arrow in $\mathbf{Set}^{r}(\mathsf{A})$, then $\mathbf{I}_{\mathsf{A}}(\mathsf{f})$ is a set-fibred arrow.
Proof.
Take the object $\mathsf{D}$ in $\mathbf{Set}^{r}(\Sigma(\mathsf{A},\mathsf{C}))$ defined as
$\mathsf{D}:=\{y\ |\ p_{1}(y)\ \varepsilon\ \mathsf{B}[p_{1}(x)/x]\ \wedge p_{2}(y)=p_{2}(x)\wedge p_{2}(y)=\{\mathbf{n}_{\mathsf{f}}\}(p_{1}(x),p_{1}(y))\}$
then $\mathbf{I}_{\mathsf{A}}(\mathsf{f})$, seen as an object in $\bm{\mathcal{C}}_{r}/(\Sigma(\mathsf{A},\mathsf{C})$, is isomorphic to $\mathbf{I}_{\Sigma(\mathsf{A},\mathsf{C})}(\mathsf{D})$. ∎
Using the encoding of $\Sigma$ and $\Pi$ we obtain the following properties about substitution along set-fibred maps.
Lemma 3.9.
If $\mathsf{f}:\mathsf{A}\to\mathsf{B}$ is set-fibred then $\mathbf{Set}^{r}_{\mathsf{f}}$ has a left adjoint.
Proof.
Consider the arrow $\mathsf{p}_{1}^{\Sigma}=[\mathbf{p}_{1}]_{\approx}:\Sigma(\mathsf{B},\tau_{\mathsf{B}}(\mathbf{n}))\rightarrow\mathsf{B}$. Then
$\Sigma_{\mathsf{p}_{1}^{\Sigma}}(\tau_{\Sigma(\mathsf{B},\tau_{\mathsf{B}}(\mathbf{n}))}(\mathbf{m}))\simeq\tau_{\mathsf{B}}(\Lambda x.\{\bm{\sigma}\}(\{\mathbf{n}\}(x),\Lambda y.\{\mathbf{m}\}(\{\mathbf{pair}\}(x,y))))$
where $\bm{\sigma}$ is a numeral representing the construction of codes of $\Sigma$-types in the universe of realized sets [IMMS18]. ∎
Lemma 3.10.
If $\mathsf{f}:\mathsf{A}\rightarrow\mathsf{B}$ in $\bm{\mathcal{C}}_{r}$ is set-fibred, for every object $\mathsf{C}$ in $\mathbf{Set}^{r}(\mathsf{A})$ the object $\Pi^{r}_{\mathsf{f}}(\mathsf{C})$ is isomorphic in $\mathbf{Col}^{r}(\mathsf{B})$ to an object of $\mathbf{Set}^{r}(\mathsf{B})$.
Proof.
Consider the arrow $\mathsf{p}_{1}^{\Sigma}=[\mathbf{p}_{1}]_{\approx}:\Sigma(\mathsf{B},\tau_{\mathsf{B}}(\mathbf{n}))\rightarrow\mathsf{B}$. Then, considering the definition of $\Pi^{r}_{\mathsf{f}}$ in [MM21, Theorem 4.9], we get that
$\Pi_{\mathsf{p}_{1}^{\Sigma}}(\tau_{\Sigma(\mathsf{B},\tau_{\mathsf{B}}(\mathbf{n}))}(\mathbf{m}))\simeq\tau_{\mathsf{B}}(\Lambda x.\{\bm{\pi}\}(\{\mathbf{n}\}(x),\Lambda y.\{\mathbf{m}\}(\{\mathbf{pair}\}(x,y))))$
where $\bm{\pi}$ is a numeral representing the construction of codes of $\Pi$-types in the universe of realized sets [IMMS18]. ∎
Using set-fibred arrows, we obtain the following extension of [MM21, Theorem 4.24].
Theorem 3.11.
For every $\mathsf{A}$ in $\bm{\mathcal{C}}_{r}$, $\mathbf{Set}^{r}(\mathsf{A})$ is a finitely complete category with disjoint stable finite coproducts, parameterized list objects and weak exponentials. Moreover, if $\mathsf{f}:\mathsf{A}\rightarrow\mathsf{B}$ is set-fibred, then $\mathbf{Set}^{r}_{\mathsf{f}}$ has a left adjoint and for every object $\mathsf{C}$ in $\mathbf{Set}^{r}(\mathsf{A})$, $\Pi^{r}_{\mathsf{f}}(\mathsf{C})$ is isomorphic in $\mathbf{Col}^{r}(\mathsf{B})$ to an object of $\mathbf{Set}^{r}(\mathsf{B})$.
The embedding of $\mathbf{Set}^{r}$ in $\mathbf{Col}^{r}$ preserves finite limits, finite coproducts, parameterized list objects, and weak exponentials. ∎
3.5 Realized propositions.
Now we recall the notion of realized propositions on $\bm{\mathcal{C}}_{r}$.
Definition 3.12.
The indexed category $\overline{{\mathbf{Prop}^{r}}}:\bm{\mathcal{C}}_{r}^{op}\rightarrow\mathbf{Pos}$ is defined as the poset reflection of $\mathbf{Col}^{r}$. In this case we write $x^{\prime}\Vdash\mathsf{P}$ instead of $x^{\prime}\,\varepsilon\,\mathsf{P}$ for an object $\mathsf{P}$ in $\mathbf{Col}^{r}(\mathsf{A})$ seen as a representative of an object $[\mathsf{P}]$ of $\overline{{\mathbf{Prop}^{r}}}(\mathsf{A})$.
We recall the following definition:
Definition 3.13.
If $\bm{\mathcal{C}}$ is a finitely complete category, the doctrine of its weak subobjects $\mathbf{wSub}_{\bm{\mathcal{C}}}$ is the posetal reflection of the slice pseudofunctor associated with $\bm{\mathcal{C}}$.
As proved in [MM21, Theorem 4.19], realized propositions correspond to weak subobjects.
Theorem 3.14.
The functor $\overline{{\mathbf{Prop}^{r}}}$ is a first-order hyperdoctrine naturally isomorphic to the doctrine of weak subobjects $\mathbf{wSub}_{\bm{\mathcal{C}}_{r}}$ of $\bm{\mathcal{C}}_{r}$. ∎
If $\mathsf{f}:\mathsf{A}\to\mathsf{B}$ is an arrow in $\bm{\mathcal{C}}_{r}$, the left and right adjoint to $\overline{{\mathbf{Prop}^{r}}}_{\mathsf{f}}$ will be denoted with $\exists_{\mathsf{f}}$ and $\forall_{\mathsf{f}}$. respectively.
3.6 Realized small propositions.
Definition 3.15.
The functor $\overline{{\mathbf{Prop}^{r}_{\mathbf{s}}}}:\bm{\mathcal{C}}_{r}^{op}\rightarrow\mathbf{Pos}$ is defined as the poset reflection of the indexed category $\mathbf{Set}^{r}$. Then, fibre objects of $\overline{{{\mathbf{Prop}^{r}_{\mathbf{s}}}}}$ are called small realized propositions [^4] Note that here we use the same notation of [MM21] for the indexed categories $\mathbf{Set}^{r}$ and $\overline{{{\mathbf{Prop}^{r}_{\mathbf{s}}}}}$, that were defined in loc. cit. on the subcategory $\mathcal{S}_{r}$ of $\bm{\mathcal{C}}_{r}$ of realized sets..
We now obtain the following theorem which is a straightforward extension of [MM21, Theorem 4.27].
Theorem 3.16.
For every object $\mathsf{A}$ in $\bm{\mathcal{C}}_{r}$, $\overline{{\mathbf{Prop}^{r}_{\mathbf{s}}}}(\mathsf{A})$ is a Heyting algebra and for every arrow $\mathsf{f}$ in $\bm{\mathcal{C}}_{r}$, $(\overline{{\mathbf{Prop}^{r}_{\mathbf{s}}}})_{\mathsf{f}}$ is a homomorphism of Heyting algebras.
Moreover, if $\mathsf{f}:\mathsf{A}\rightarrow\mathsf{B}$ in $\bm{\mathcal{C}}_{r}$ is set-fibred, then $(\overline{{\mathbf{Prop}^{r}_{\mathbf{s}}}})_{\mathsf{f}}$ has left $\exists_{\mathsf{f}}$ and right $\forall_{\mathsf{f}}$ adjoints satisfying Beck-Chevalley conditions. ∎
3.7 Further notation.
For every $\mathsf{f}:=[\mathbf{n}]_{\approx}:\mathsf{A}\rightarrow\mathsf{B}$ in $\bm{\mathcal{C}}_{r}$ and every $\mathsf{C}$
in $\mathbf{Col}^{r}(\mathsf{B})$, we denote with $\Sigma(\mathsf{f},\mathsf{C})$ the arrow $[\Lambda x.\{\mathbf{pair}\}(\{\mathbf{n}\}(\{\mathbf{p}_{1}\}(x)),\{\mathbf{p}_{2}\}(x))\,]_{\approx}$ from $\Sigma(\mathsf{A},\mathbf{Col}_{\mathsf{f}}^{r}(\mathsf{C}))$ to $\Sigma(\mathsf{B},\mathsf{C})$. This arrow makes the following diagram a pullback in $\bm{\mathcal{C}}_{r}$
${{\Sigma(\mathsf{A},\mathbf{Col}_{\mathsf{f}}^{r}(\mathsf{C}))}}$${{\Sigma(\mathsf{B},\mathsf{C})}}$${{\mathsf{A}}}$${{\mathsf{B}}}$$\scriptstyle{\Sigma(\mathsf{f},\mathsf{C})}$$\scriptstyle{\mathsf{p}_{1}^{\Sigma}}$$\scriptstyle{\mathsf{p}_{1}^{\Sigma}}$$\scriptstyle{\mathsf{f}}$
for details see [MM21, Lemma 4.11].
Now let $\mathsf{A}$ be an object in $\bm{\mathcal{C}}_{r}$ and let $\mathsf{B}$ be an object of $\mathbf{Col}^{r}(\mathsf{A})$. If $\mathsf{C}$ is an object in $\mathbf{Col}^{r}(\Sigma(\mathsf{A},\mathsf{B}))$, then we define $\Sigma^{\mathsf{A}}(\mathsf{B},\mathsf{C})$ as the object in $\mathbf{Col}^{r}(\mathsf{A})$ given by $\Sigma_{\mathsf{p}^{\Sigma}_{1}}(\mathsf{C})$.
Moreover, for every arrow $\mathsf{f}:\mathsf{B}^{\prime}\to\mathsf{B}$ in $\textbf{Col}^{r}(\mathsf{A})$, we define $\Sigma^{\mathsf{A}}(\mathsf{f},\mathsf{C})$ as the arrow
$\Lambda x.\Lambda x^{\prime}.\{\mathbf{pair}\}(\{\mathbf{pair}\}(x,\{\mathbf{n}_{\mathsf{f}}\}(x,\{\mathbf{p_{2}}\}(\{\mathbf{p}_{1}\}(x^{\prime}))),\{\mathbf{p_{2}}\}(x^{\prime}))$
which makes the following diagram a pullback in $\mathbf{Col}^{r}(\mathsf{A})$
${{\Sigma^{\mathsf{A}}(\mathsf{B}^{\prime},\mathbf{Col}_{\mathsf{f}}^{r}(\mathsf{C}))}}$${{\Sigma^{\mathsf{A}}(\mathsf{B},\mathsf{C})}}$${{\mathsf{B}^{\prime}}}$${{\mathsf{B}}}$$\scriptstyle{\Sigma^{\mathsf{A}}(\mathsf{f},\mathsf{C})}$$\scriptstyle{\mathsf{p}_{1}^{\mathsf{A},\Sigma}}$$\scriptstyle{\mathsf{p}_{1}^{\mathsf{A},\Sigma}}$$\scriptstyle{\mathsf{f}}$
where $\mathsf{p}_{1}^{\mathsf{A},\Sigma}:=[\Lambda x.\Lambda x^{\prime}.\{\mathbf{p_{2}}\}(\{\mathbf{p_{1}}\}(x^{\prime}))]_{\approx}$.
We now define a currying isomorphism in $\bm{\mathcal{C}}_{r}$ for every object $\mathsf{A}\in\bm{\mathcal{C}}_{r}$, $\mathsf{B}\in\mathbf{Col}^{r}(\mathsf{A})$ and $\mathsf{C}\in\mathbf{Col}^{r}(\Sigma(\mathsf{A},\mathsf{B}))$
$\mathsf{cur}:\Sigma(\Sigma(\mathsf{A},\mathsf{B}),\mathsf{C})\to\Sigma(\mathsf{A},\Sigma^{\mathsf{A}}(\mathsf{B},\mathsf{C}))$
given by
$\Lambda x.\{\mathbf{pair}\}(\{\mathbf{p}_{1}\}(\{\mathbf{p}_{1}\}(x)),x).$ (8)
4 An internal language for ${\mathbf{Prop}^{r}}$ and $\mathbf{Col}^{r}$
As done in [vO08], we consider the internal language of the functors ${\mathbf{Prop}^{r}}$ and $\mathbf{Col}^{r}$. The former is in the form of a first-order typed language with equality, and the latter is in the form of one with dependent types.
A context $\Gamma$ is an expression of the form $[x_{0}:\mathsf{A}_{0},\dots,x_{n}:\mathsf{A}_{n}(x_{1},\dots,x_{n-1})]$. A substitution of contexts $\bar{t}:\Gamma\to\Delta$, where $\Delta:=[y_{0}:\mathsf{B}_{0},\dots,y_{m}:\mathsf{B}_{m}(y_{0},\dots,y_{m-1})]$ is a list of terms $\bar{t}:=[t_{0},\dots,t_{m}]$ such that $t_{i}:\mathsf{B}_{i}(t_{0},\dots,t_{i-1})$ is in context $[\Gamma]$ for $0\leq i\leq m$.
We will adopt the notation $\mathsf{C}_{x}(y)$ for a type in context $x:\mathsf{A},y:\mathsf{B}(x)$ and $\mathsf{f}_{x}:\mathsf{B}(x)\to\mathsf{B}^{\prime}(x)$ for a term of type $\mathsf{B}^{\prime}(x)$ in context $x:\mathsf{A},y:\mathsf{B}(x)$.
Recall from [MM21], that contexts $\Gamma$ are interpreted as objects $\|\Gamma\|$ of the category $\bm{\mathcal{C}}_{r}$ and substitutions of contexts as arrows of $\bm{\mathcal{C}}_{r}$. A formula $\psi[\Gamma]$ in context $\Gamma$ is interpreted as an element of ${\mathbf{Prop}^{r}}(\|\Gamma\|)$, while a dependent type $\mathsf{A}[\Gamma]$ in context $\Gamma$ is interpreted as an object of $\mathbf{Col}^{r}(\|\Gamma\|)$. Given the interpretations $\|\bar{t}\|:\|\Gamma\|\to\|\Delta\|$ and $\|\psi[\Delta]\|\in{\mathbf{Prop}^{r}}(\|\Delta\|)$, or $\|\mathsf{A}[\Delta]\|\in\mathbf{Col}^{r}(\|\Delta\|)$, then the interpretation of $\psi(\bar{t})[\Gamma]$, or $\mathsf{A}(\bar{t})[\Gamma]$, is respectively given by ${\mathbf{Prop}^{r}}_{\|t\|}(\|\psi\|)$, or $\mathbf{Col}^{r}_{\|t\|}(\|\mathsf{A}\|)$.
Furthermore, we assume that given an interpretation of $\mathsf{A}[\Gamma]$ as an object $\|\mathsf{A}\|\in\mathbf{Col}^{r}(\|\Gamma\|)$, then the interpretation of the context $[\Gamma,x:\mathsf{A}]$ is given by the object $\Sigma(\|\Gamma\|,\|\mathsf{A}\|)$ of $\bm{\mathcal{C}}_{r}$. In particular, given a substitution $\bar{t}:\Gamma\to\Delta$ and a type in context $\mathsf{A}[\Delta]$, then the substitution $[\bar{t},y]:[\Gamma,y:\mathsf{A}(\bar{t})]\to[\Delta,z:\mathsf{A}]$ is interpreted as the arrow
$\Sigma(\|\bar{t}\|,\|\mathsf{A}\|):\Sigma(\|\Gamma\|,\mathbf{Col}^{r}_{\|\bar{t}\|}(\|\mathsf{A}\|))\to\Sigma(\|\Delta\|,\|\mathsf{A}\|)$
If $\mathsf{A}[\Gamma]$ and $\mathsf{B}[\Gamma]$ are types in the same context and $f:\mathsf{A}\to\mathsf{B}$ is a term in context $[\Gamma,x:\mathsf{A}]$, then $f$ is interpreted as an arrow $\|f\|:\|\mathsf{A}\|\to\|\mathsf{B}\|$ in $\mathbf{Col}^{r}(\|\Gamma\|)$ and the substitution $[id_{\Gamma},f]:[\Gamma,x:\mathsf{A}]\to[\Gamma,y:\mathsf{B}]$ is interpreted as the arrow
$\mathbf{I}_{\|\Gamma\|}(\|f\|):\Sigma(\|\Gamma\|,\|\mathsf{A}\|)\to\Sigma(\|\Gamma\|,\|\mathsf{B}\|)$
If $\psi[\Gamma]$ is a formula in context, we define its validity ${\mathbf{Prop}^{r}}\vdash\psi[\Gamma]$ if and only if $\top_{\|\Gamma\|}\leq_{\|\Gamma\|}\|\psi[\Gamma]\|$. Similarly, for a type $\mathsf{A}[\Gamma]$ in context, we define $\mathbf{Col}^{r}\vdash\mathsf{A}[\Gamma]$ if and only if there exists an arrow $1_{\|\Gamma\|}\to\|\mathsf{A}[\Gamma]\|$.
[…]
5 The predicative effective topos $\mathbf{pEff}$
In [MM21], the predicative effective topos $\mathbf{pEff}$ is defined as the base category $\mathcal{Q}_{\overline{\mathbf{Prop}^{r}}}$ of the elementary quotient completion of the elementary doctrine $\overline{\mathbf{Prop}^{r}}$. It is the category whose objects are the pairs $(\mathsf{A},[\mathsf{R}])$ where $\mathsf{A}$ is an object of the category $\bm{\mathcal{C}}_{r}$ and $[\mathsf{R}]$ is an object of $\overline{\mathbf{Prop}^{r}}(\mathsf{A}\times\mathsf{A})$ such that
  1. 1.
    $\top_{\mathsf{A}}\leq_{\mathsf{A}}\overline{\mathbf{Prop}^{r}}_{\Delta_{\mathsf{A}}}([\mathsf{R}])$;
  2. 2.
    $[\mathsf{R}]\leq_{\mathsf{A}\times\mathsf{A}}\overline{\mathbf{Prop}^{r}}_{\langle p_{2},p_{1}\rangle}([\mathsf{R}])$, where $p_{1},p_{2}:\mathsf{A}\times\mathsf{A}\to\mathsf{A}$ are the projections;
  3. 3.
    $\overline{\mathbf{Prop}^{r}}_{\langle p_{1},p_{2}\rangle}([\mathsf{R}])\wedge\overline{\mathbf{Prop}^{r}}_{\langle p_{2},p_{3}\rangle}([\mathsf{R}])\leq_{\mathsf{A}\times\mathsf{A}\times\mathsf{A}}\overline{\mathbf{Prop}^{r}}_{\langle p_{1},p_{3}\rangle}([\mathsf{R}])$, where $p_{1},p_{2},p_{3}:\mathsf{A}\times\mathsf{A}\times\mathsf{A}\to\mathsf{A}$ are the projections;
Equivalently, the above conditions state that $[\mathsf{R}]$ is an equivalence relation for the internal logic of the doctrine $\overline{\mathbf{Prop}^{r}}$, see [MM21].
An arrow from $(\mathsf{A},[\mathsf{R}])$ to $(\mathsf{B},[\mathsf{S}])$ is an equivalence class $[\mathsf{f}]$ of arrows $\mathsf{f}:\mathsf{A}\rightarrow\mathsf{B}$ in $\bm{\mathcal{C}}_{r}$ such that $[\mathsf{R}]\leq_{\mathsf{A}\times\mathsf{A}}\overline{\mathbf{Prop}^{r}}_{\mathsf{f}\times\mathsf{f}}([\mathsf{S}])$, where $\mathsf{f}\sim\mathsf{g}$ if and only if $\top_{A}\leq_{\mathsf{A}}\overline{\mathbf{Prop}^{r}}_{\langle\mathsf{f},\mathsf{g}\rangle}([\mathsf{S}])$.
Moreover, we recall from [MM21] that $\mathbf{pEff}$ is equivalent to the exact completion $(\bm{\mathcal{C}}_{r})_{ex/lex}$ of the left exact category $\bm{\mathcal{C}}_{r}$. Indeed, this follows from theorem 3.14 and the fact that the elementary quotient completion of a weak subobject doctrine on a lex category $C$ is equivalent to the exact on lex completion of $C$ (see [MR13b, MR13a]).
The following appears as [MM21, Theorem 5.5].
Theorem 5.1.
The category $\mathbf{pEff}$ is a locally cartesian closed list arithmetic pretopos.
6 Families of collections and sets over $\mathbf{pEff}$
[…]
The intuitive idea behind the definition of a family of collections over $(\mathsf{A},\mathsf{R})$ is the following:
  • first we take a family of realized collections $\mathsf{B}(a)$ over $\mathsf{A}$ in $\bm{\mathcal{C}}_{r}$;
  • then, we take an equivalence relation $\mathsf{S}_{a}(b,b^{\prime})$ on each $\mathsf{B}(a)$ (depending on $\mathsf{A}$);
  • furthermore, we take a family of transport maps $b\mapsto\sigma_{a,a^{\prime}}(r,b)$ for $a,a^{\prime}\,\varepsilon\,\mathsf{A}$, $b\,\varepsilon\,B(a)$ and $r\Vdash\mathsf{R}(a,a^{\prime})$ such that $\sigma_{a,a^{\prime}}(r,b)\,\varepsilon\,\mathsf{B}(a^{\prime})$;
  • Finally, we ask $\sigma$ to satisfy the following conditions, which makes it a sort of action:
    1. 1.
      if $S_{a}(b,b^{\prime})$ and $r\Vdash\mathsf{R}(a,a^{\prime})$, then $\mathsf{S}_{a^{\prime}}(\sigma_{a,a^{\prime}}(r,b),\sigma_{a,a^{\prime}}(r,b^{\prime}))$, that is $\sigma$ respects $\mathsf{S}$;
    2. 2.
      if $r\Vdash\mathsf{R}(a,a)$ and $b\,\varepsilon\,\mathsf{B}(a)$, then $\mathsf{S}_{a}(b,\sigma_{a,a}(r,b)))$, that is $\sigma$ is well-behaved with respect to identities;
    3. 3.
      if $r\Vdash\mathsf{R}(a,a^{\prime})$, $r^{\prime}\Vdash\mathsf{R}(a^{\prime},a^{\prime\prime})$, $r^{\prime\prime}\Vdash\mathsf{R}(a,a^{\prime\prime})$ and $b\,\varepsilon\,\mathsf{B}(a)$, then
      $\mathsf{S}_{a^{\prime\prime}}(\sigma_{a,a^{\prime\prime}}(r^{\prime\prime},b),\sigma_{a^{\prime},a^{\prime\prime}}(r^{\prime},\sigma_{a,a^{\prime}}(r,b))))$
      that is $\sigma$ is well-behaved with respect to composition.
    Observe that, taking $a^{\prime}=a^{\prime\prime}$ and a witness for the reflexivity of $\mathsf{R}$, conditions 2. and 3. above imply the following condition stating that $\sigma$, once we fix two related objects $a,a^{\prime}$, does not depend on the particular realizer of $\mathsf{R}(a,a^{\prime})$:
    1. 4.
      ${\mathsf{S}}_{a^{\prime}}(\sigma_{a,a^{\prime}}(r,b),\sigma_{a,a^{\prime}}(r^{\prime},b))$ for every $b\,\varepsilon\,\mathsf{B}(a)$, $a,a^{\prime}\varepsilon\mathsf{A}$, $r,r^{\prime}\Vdash\mathsf{R}(a,a^{\prime})$;
To formulate this in a precise way this idea we need first to introduce the notion of fibered ${\overline{\textbf{Prop}^{r}}}$-equivalence relation:
Definition 6.1.
Let $\mathsf{A}$ be an object of $\bm{\mathcal{C}}_{r}$ and let $\mathsf{B}$ be an object of $\textbf{Col}^{r}(\mathsf{A})$. A fibered ${\overline{\mathbf{Prop}}^{r}}$-equivalence relation on $\mathsf{B}$, or simply a fibered equivalence relation on $\mathsf{B}$, is an object $[\mathsf{S}]$ in ${\overline{\textbf{Prop}^{r}}}(\Sigma(\mathsf{A},\mathsf{B}\times\mathsf{B}))$ such that
  1. 1.
    $\top_{\Sigma(\mathsf{A},\mathsf{B})}\leq_{\Sigma(\mathsf{A},\mathsf{B})}{\overline{\textbf{Prop}^{r}}}_{\textbf{I}_{\mathsf{A}}(\Delta_{\mathsf{B}})}([\mathsf{S}])$
  2. 2.
    $[\mathsf{S}]\leq_{\Sigma(\mathsf{A},\mathsf{B}\times\mathsf{B})}{\overline{\textbf{Prop}^{r}}}_{\textbf{I}_{\mathsf{A}}(\langle p_{2},p_{1}\rangle)}([\mathsf{S}])$
  3. 3.
    ${\overline{\textbf{Prop}^{r}}}_{\textbf{I}_{\mathsf{A}}(\langle p_{1},p_{2}\rangle)}([\mathsf{S}])\wedge{\overline{\textbf{Prop}^{r}}}_{\textbf{I}_{\mathsf{A}}(\langle p_{2},p_{3}\rangle)}([\mathsf{S}])\leq_{\Sigma(\mathsf{A},\mathsf{B}\times\mathsf{B}\times\mathsf{B})}{\overline{\textbf{Prop}^{r}}}_{\textbf{I}_{\mathsf{A}}(\langle p_{1},p_{3}\rangle)}([\mathsf{S}])$.
Definition 6.2.
Let $(\mathsf{\mathsf{A}},[\mathsf{R}])$ be an object of pEff. An extensional dependent collection on $(\mathsf{\mathsf{A}},[\mathsf{R}])$, with respect to ${\mathsf{R}}$, is a triple $(\mathsf{B},[\mathsf{S}],\sigma)$ where $\mathsf{B}\in\textbf{Col}^{r}(\mathsf{A})$, $[\mathsf{S}]$ is a dependent ${\overline{\textbf{Prop}^{r}}}$-equivalence relation on $\mathsf{B}$ and
$\sigma:\textbf{Col}^{r}_{p_{1}}(\mathsf{B})\times\mathsf{R}\to\textbf{Col}^{r}_{p_{2}}(\mathsf{B})$
is an arrow in $\textbf{Col}^{r}(\mathsf{A}\times\mathsf{A})$ satisfying
  1. 1.
    $\sigma$ preserves $[\mathsf{S}]$ i.e. for the arrow
    ${{\Sigma(\mathsf{A}\times\mathsf{A},\textbf{Col}^{r}_{p_{1}}(\mathsf{B}\times\mathsf{B})\times\mathsf{R})}}$${{\Sigma(\mathsf{A}\times\mathsf{A},\textbf{Col}^{r}_{p_{2}}(\mathsf{B}\times\mathsf{B}))}}$$\scriptstyle{\textbf{I}((\sigma\circ\langle p_{1},p_{3}\rangle)\times(\sigma\circ\langle p_{2},p_{3}\rangle))}$
    it holds
    $\overline{\textbf{Prop}^{r}}_{{\Sigma(p_{1},\mathsf{B}\times\mathsf{B})\circ\textbf{I}(p_{1})}}[\mathsf{S}]\leq\overline{\textbf{Prop}^{r}}_{\Sigma(p_{2},\mathsf{B}\times\mathsf{B})\circ{\textbf{I}((\sigma\circ\langle p_{1},p_{3}\rangle)\times(\sigma\circ\langle p_{2},p_{3}\rangle))}}[\mathsf{S}]$
    where $\Sigma(p_{i},\mathsf{B}\times\mathsf{B}):\Sigma(\mathsf{A}\times\mathsf{A},\textbf{Col}^{r}_{p_{1}}(\mathsf{B}\times\mathsf{B}))\to{\Sigma(\mathsf{A},\mathsf{B}\times\mathsf{B})}$, for $i=1,2$.
  2. 2.
    $\sigma$ preserves the identities, i.e. for the unique arrow $\rho$ making the following diagram commute
    ${{\Sigma(\mathsf{A},\mathsf{B})}}$${{\Sigma(\mathsf{A},\mathsf{B}\times\textbf{Col}^{r}_{\Delta}(\mathsf{R}))}}$${{\Sigma(\mathsf{A},\mathsf{B}\times\mathsf{B})}}$${{\Sigma(\mathsf{A},\mathsf{B})}}$$\scriptstyle{\textbf{I}(p_{1})}$$\scriptstyle{\textbf{I}(\textbf{Col}^{r}_{\Delta}\sigma)}$$\scriptstyle{\textbf{I}(p_{1})}$$\scriptstyle{\textbf{I}(p_{2})}$$\scriptstyle{\rho}$
    then $\top\leq{\overline{\textbf{Prop}^{r}}}_{\rho}([\mathsf{S}])$.
  3. 3.
    $\sigma$ preserves the composition, i.e. for the unique arrow $\tau$  [^5] Here we use the notation $\mathsf{A}^{3}:=\mathsf{A}\times\mathsf{A}\times\mathsf{A}$
    ${{\Sigma(\mathsf{A}^{3},\textbf{Col}^{r}_{p_{1}}(\mathsf{B})\times\textbf{Col}^{r}_{\langle p_{1},p_{2}\rangle}(\mathsf{R})\times\textbf{Col}^{r}_{\langle p_{2},p_{3}\rangle}(\mathsf{R})\times\textbf{Col}^{r}_{\langle p_{1},p_{3}\rangle}(\mathsf{R}))}}$${\Sigma(\mathsf{A}^{3},\textbf{Col}^{r}_{p_{3}}(\mathsf{B}\times\mathsf{B}))}$$\scriptstyle{\tau}$
    such that
    • $\mathbf{I}(p_{2})\circ\tau=\mathbf{I}(\mathbf{Col}^{r}_{\langle p_{1},p_{3}\rangle}(\sigma)\circ\langle p_{1},p_{4}\rangle)$
    • $\mathbf{I}(p_{1})\circ\tau=\mathbf{I}(p_{1})\circ\mathbf{I}(\mathbf{Col}^{r}_{\langle p_{2},p_{3}\rangle}(\sigma)\circ(\mathbf{Col}^{r}_{\langle p_{1},p_{2}\rangle}(\sigma)\times id)\times id)$
    it holds that $\top\leq{\overline{\textbf{Prop}^{r}}}_{\Sigma(p_{3},\mathsf{B}\times\mathsf{B})\circ\tau}([\mathsf{S}])$.
Definition 6.3.
Let $(\mathsf{A},[\mathsf{R}])$ be an object in $\mathbf{pEff}$. A morphism between extensional dependent collections over $(\mathsf{A},[\mathsf{R}])$ w.r.t. ${\mathsf{R}}$ from $(\mathsf{B},[\mathsf{S}],\sigma)$ to $(\mathsf{C},[\mathsf{H}],\nu)$
is an equivalence class ${[\mathsf{f}]}_{\cong}$ of arrows $\mathsf{f}:\mathsf{B}\to\mathsf{C}$ in $\textbf{Col}^{r}(\mathsf{A})$ such that
  1. 1.
    $\mathsf{f}$ respects the equivalence relations:
    $[\mathsf{S}]\leq\overline{\textbf{Prop}^{r}}_{\textbf{I}(\mathsf{f}\times\mathsf{f})}([\mathsf{H}])$
  2. 2.
    $\mathsf{f}$ commutes with the transport maps $\mathsf{H}(\nu_{a,a^{\prime}}(r,\mathsf{f}(b)),\mathsf{f}(\sigma_{a,a^{\prime}}(r,b)))$:
    $\top\leq\overline{\textbf{Prop}^{r}}_{\Sigma(p_{2},\mathsf{C}\times\mathsf{C})\circ\mathsf{squ}}([\mathsf{H}])$
    where $\mathsf{squ}:=\textbf{I}(\langle\nu\circ(\textbf{Col}^{r}_{p_{1}}(\mathsf{f})\times 1_{\mathsf{R}}),\textbf{Col}^{r}_{p_{2}}(\mathsf{f})\circ\sigma\rangle)$;
and $\mathsf{f}\sim\mathsf{g}$ if and only if $[\mathsf{S}]\leq\overline{\textbf{Prop}^{r}}_{\textbf{I}(\mathsf{f}\times\mathsf{g})}([\mathsf{H}])$.
Remark 6.4.
The above notion was inspired by that of extensional dependent collection in the context of dependent type theory in [Mai09].
Definition 6.5.
Let $(\mathsf{\mathsf{A}},[\mathsf{R}])$ be an object of pEff. An extensional dependent set on $(\mathsf{\mathsf{A}},[\mathsf{R}])$ w.r.t ${\mathsf{R}}$ is an extensional dependent collection $(\mathsf{B},[\mathsf{S}],\sigma)$ on $(\mathsf{\mathsf{A}},[\mathsf{R}])$ w.r.t ${\mathsf{R}}$ such that $\mathsf{B}\in\textbf{Set}^{r}(\mathsf{A})$ and $[\mathsf{S}]\in{\overline{\textbf{Prop}_{s}^{r}}}(\Sigma(\mathsf{A},\mathsf{B}\times\mathsf{B}))$.
Definition 6.6.
Given $(\mathsf{A},[\mathsf{R}])\in\mathbf{pEff}$, we denote with $Dep_{\mathbf{pEff}}(\mathsf{A},{\mathsf{R}})$ the category of extensional dependent collections on $(\mathsf{\mathsf{A}},[\mathsf{R}])$ w.r.t ${\mathsf{R}}$ and arrows between them. Furthermore, we denote with $Dep_{\mathbf{pEff}_{set}}(\mathsf{A},{\mathsf{R}})$ its full subcategory on extensional dependent sets.
Lemma 6.7.
Let $(\mathsf{A},[\mathsf{R}])$ be an object in $\mathbf{pEff}$ and let $\mathsf{R},\mathsf{R}^{\prime}$ be two representatives of $[\mathsf{R}]$. The categories $Dep_{\mathbf{pEff}}(\mathsf{A},{\mathsf{R}})$ and $Dep_{\mathbf{pEff}}(\mathsf{A},{\mathsf{R}}^{\prime})$ are equivalent. The same holds for extensional dependent sets.
Proof.
The fact that $\mathsf{R}$ and $\mathsf{R}^{\prime}$ are in $[\mathsf{R}]\in\overline{{\mathbf{Prop}^{r}}}(\mathsf{A}\times\mathsf{A})$ imply that there exist two arrows $h:\mathsf{R}\to\mathsf{R}^{\prime}$ and $k:\mathsf{R}^{\prime}\to\mathsf{R}$ in $\mathbf{Col}^{r}(\mathsf{A}\times\mathsf{A})$. We can construct the functor
$-\circ k:Dep_{\mathbf{pEff}}(\mathsf{A},{\mathsf{R}})\to Dep_{\mathbf{pEff}}(\mathsf{A},{\mathsf{R}}^{\prime})$
that sends an extensional dependent collection $(\mathsf{B},[\mathsf{S}],\sigma)$ to the triple $(\mathsf{B},[\mathsf{S}],\sigma\circ(id\times k))$ and every arrow into itself. It is straightforward to prove this functor is an equivalence considering the similar functor
$-\circ h:Dep_{\mathbf{pEff}}(\mathsf{A},{\mathsf{R}}^{\prime})\to Dep_{\mathbf{pEff}}(\mathsf{A},{\mathsf{R}}).$
In the following proposition, we show directly that extensional dependent collections induce objects in the slice categories of $\mathbf{pEff}$.
Proposition 6.8.
Let $(\mathsf{A},[\mathsf{R}])$ be an object in $\mathbf{pEff}$. There exists a full and faithful essentially surjective functor
$\mathbf{K}:Dep_{\mathbf{pEff}}(\mathsf{A},{\mathsf{R}})\to\mathbf{pEff}/(\mathsf{A},{[\mathsf{R}]}).$
Proof.
For every object $(\mathsf{B},\mathsf{S},\sigma)$ we define
$\mathbf{K}(\mathsf{B},[\mathsf{S}],\sigma):=[\mathsf{p}_{1}^{\Sigma}]_{\cong}:(\Sigma(\mathsf{A},\mathsf{B}),\exists_{d_{{\mathsf{R}}}}(\overline{\mathbf{Prop}^{r}}_{t_{\sigma}}([\mathsf{S}]))\to(\mathsf{A},[\mathsf{R}])$ (9)
for the commutative diagram
${{\Sigma(\mathsf{A}\times\mathsf{A},\mathsf{R}\times\mathbf{Col}^{r}_{p_{1}}(\mathsf{B})\times\mathbf{Col}^{r}_{p_{2}}(\mathsf{B}))}}$${{\Sigma(\mathsf{A}\times\mathsf{A},\mathsf{R}\times\mathbf{Col}^{r}_{p_{1}}(\mathsf{B})\times\mathbf{Col}^{r}_{p_{2}}(\mathsf{B}))}}$${{\Sigma(\mathsf{A}\times\mathsf{A},\mathbf{Col}^{r}_{p_{2}}(\mathsf{B})\times\mathbf{Col}^{r}_{p_{2}}(\mathsf{B}))}}$${{\Sigma(\mathsf{A}\times\mathsf{A},\mathbf{Col}^{r}_{p_{1}}(\mathsf{B})\times\mathbf{Col}^{r}_{p_{2}}(\mathsf{B}))}}$${{\Sigma(\mathsf{A},\mathsf{B}\times\mathsf{B})}}$${{\Sigma(\mathsf{A},\mathsf{B})\times\Sigma(\mathsf{A},\mathsf{B})}}$$\scriptstyle{\mathbf{I}(\langle\sigma\circ\langle p_{1},p_{2}\rangle,p_{3}\rangle)}$$\scriptstyle{\Sigma(p_{2},\mathsf{B}\times\mathsf{B})}$$\scriptstyle{k}$$\scriptstyle{t_{\sigma}}$$\scriptstyle{\mathbf{I}(\langle p_{2},p_{3}\rangle)}$$\scriptstyle{d_{{\mathsf{R}}}}$
Where $k$ is the isomorphism given by $\langle\Sigma(p_{1},\mathsf{B})\circ\mathbf{I}(p_{1}),\Sigma(p_{2},\mathsf{B})\circ\mathbf{I}(p_{2})\rangle$.
Informally, the domain of the arrow above has as elements those pairs $(a,b)$ with $a\in\mathsf{A}$ and $b\in\mathsf{B}(a)$. A realizer for an equivalence of two such pairs $(a,b)$ and $(a^{\prime},b^{\prime})$ is a pair $(r,s)$ such that $r\Vdash\mathsf{R}(a,a^{\prime})$ and $s\Vdash\mathsf{S}_{a^{\prime}}(\sigma_{a,a^{\prime}}(r,b),b^{\prime})$.
For $[\mathsf{f}]:(\mathsf{B},[\mathsf{S}],\sigma)\rightarrow(\mathsf{C},[\mathsf{H}],\eta)$ we define $\mathbf{K}([\mathsf{f}])$ as $[\mathbf{I}(\mathsf{f})]$. These assignments will define a functor.
Since every arrow $h$ in $\bm{\mathcal{C}}_{r}$ from $\Sigma(\mathsf{A},\mathsf{B})$ to $\Sigma(\mathsf{A},\mathsf{C})$ such that $\mathsf{p}_{1}^{\Sigma}\circ h=\mathsf{p}_{1}^{\Sigma}$ has the form $\mathbf{I}(\widetilde{h})$ for some $\widetilde{h}:\mathsf{B}\rightarrow\mathsf{C}$ in $\mathbf{Col}^{r}(\mathsf{A})$, one obtains that every arrow $h$ in $\mathbf{pEff}/(\mathsf{A},[\mathsf{R}])$ between objects in the image of $\mathbf{K}$ has the form $[\mathbf{I}(\widetilde{h})]$. The fullness of $\mathbf{K}$ follows observing that such an arrow $\widetilde{h}$ determines in fact an arrow in $Dep_{\mathbf{pEff}}(\mathsf{A},\mathsf{R})$. This can be easily obtained using the properties of the original arrow $h$.
Faithfulness is immediate once one notice that equivalence of arrows $h$’s in $\mathbf{pEff}/(\mathsf{A},[\mathsf{R}])$ in the image of $\mathbf{K}$ corresponds exactly to equivalence of the corresponding $\widetilde{h}$’s in $Dep_{\mathbf{pEff}}(\mathsf{A},\mathsf{R})$.
Now given an object $[\mathsf{f}]_{\cong}:(\mathsf{A}^{\prime},[\mathsf{R}^{\prime}])\to(\mathsf{A},[\mathsf{R}])$ in $\mathbf{pEff}/(\mathsf{A},\mathsf{R})$, we build a triple $(\mathsf{B}^{\mathsf{f}},[\mathsf{S}]^{\mathsf{f}},\sigma^{\mathsf{f}})$ in $Dep_{\mathbf{pEff}}(\mathsf{A},{\mathsf{R}})$. In order to do that, we consider the arrow $p_{1}:\mathsf{A}\times\mathsf{A}^{\prime}\to\mathsf{A}$ and define $\mathsf{B}^{\mathsf{f}}:=\Sigma_{p_{1}}(\mathbf{Col}^{r}_{(\mathsf{id}_{\mathsf{A}}\times\mathsf{f})}(\mathsf{R}))$. Intuitively, $\mathsf{B}^{\mathsf{f}}(a)$ is given by pairs $(a^{\prime},k)$ where $a^{\prime}\in\mathsf{A}^{\prime}$ and $k\Vdash\mathsf{R}(a,\mathsf{f}(a^{\prime}))$. Then we define
$[\mathsf{S}]^{\mathsf{f}}:=\overline{\mathbf{Prop}^{r}}_{(\mathsf{snd}^{\mathsf{f}}\times\mathsf{snd}^{\mathsf{f}})\circ\langle\mathbf{I}_{\mathsf{A}}(p_{1}),\mathbf{I}_{\mathsf{A}}(p_{2})\rangle}([\mathsf{R}^{\prime}])$
where
$\mathsf{snd}^{\mathsf{f}}:=[\Lambda x.\{\mathbf{p}_{2}\}(\{\mathbf{p}_{1}\}(\{\mathbf{p}_{2}\}(x)))]_{\approx}:\Sigma(\mathsf{A},\Sigma_{p_{1}}(\mathbf{Col}^{r}_{(\mathsf{id}_{\mathsf{A}}\times\mathsf{f})}(\mathsf{R})))\to\mathsf{A}^{\prime}$
and $\mathbf{I}_{\mathsf{A}}(p_{i}):\Sigma(\mathsf{A},\Sigma_{p_{1}}(\mathbf{Col}^{r}_{(\mathsf{id}_{\mathsf{A}}\times\mathsf{f})}(\mathsf{R})\times\mathbf{Col}^{r}_{(\mathsf{id}_{\mathsf{A}}\times\mathsf{f})}(\mathsf{R})))\to\Sigma(\mathsf{A},\Sigma_{p_{1}}(\mathbf{Col}^{r}_{(\mathsf{id}_{\mathsf{A}}\times\mathsf{f})}(\mathsf{R})))$, for $i=1,2$, to be the relation $\mathsf{R}^{\prime}$ on the $\mathsf{A}^{\prime}$ components of $\mathsf{B}^{\mathsf{f}}$.
Finally, we can define $\sigma^{\mathsf{f}}$ as the arrow which sends a triple $((a^{\prime},k),r)$ with $(a^{\prime},k)\in\mathsf{B}^{\mathsf{f}}(a_{1})$ and $r\Vdash\mathsf{R}(a_{1},a_{2})$ to the pair $(a^{\prime},\mathsf{tr}(\mathsf{sym}(r),k))\in\mathsf{B}^{\mathsf{f}}(a_{2})$ where $\mathsf{tr}(\mathsf{sym}(r),k)\Vdash\mathsf{R}(a_{2},\mathsf{f}(a^{\prime}))$ and $\mathsf{refl,sym}$ and $\mathsf{tr}$ are some chosen witnesses of the reflexivity, symmetry and transitivity of $\mathsf{R}$. We leave this to the reader.
Now we can observe that $\mathbf{K}(\mathsf{B}^{\mathsf{f}},[\mathsf{S}]^{\mathsf{f}},\sigma^{\mathsf{f}})\cong[\mathsf{f}]_{\cong}$. Indeed, there is an obvious arrow from $\Sigma(\mathsf{A},\mathsf{B}^{\mathsf{f}})$ to $\mathsf{A}^{\prime}$ which takes the $\mathsf{A}^{\prime}$ component of $\mathsf{B}$ to itself. Vice versa, given an element $a^{\prime}\in\mathsf{A}^{\prime}$, we can consider the triple $(\mathsf{f}(a^{\prime}),(a^{\prime},\mathsf{refl}))\in\Sigma(\mathsf{A},\mathsf{B})$. These arrows are inverse thanks to the relations $[\mathsf{R}^{\prime}]$ and $\exists_{d_{{\mathsf{R}}}}(\overline{\mathbf{Prop}^{r}}_{t_{\sigma}}([\mathsf{S}]^{\mathsf{f}}))$. ∎
Remark 6.9.
The above result provides an equivalence between the categories $Dep_{\mathbf{pEff}}(\mathsf{A},\mathsf{R})$ and $\mathbf{pEff}/(\mathsf{A},[\mathsf{R}])$ in case we assume the axiom of countable choice in the meta-theory.
Otherwise, we only obtain that $\mathbf{K}:Dep_{\mathbf{pEff}}(\mathsf{A},\mathsf{R})\rightarrow\mathbf{pEff}/(\mathsf{A},[\mathsf{R}])$ preserves and reflects all limits, colimits and exponentials.
Remark 6.10.
If $(\mathsf{A},[\mathsf{R}])$ is an object in $\mathbf{pEff}$ and $\mathsf{R},\mathsf{R}^{\prime}$ are two representatives of $[\mathsf{R}]$ as in Lemma 6.7, then the image through $\mathbf{K}$ of the categories $Dep_{\mathbf{pEff}}(\mathsf{A},{\mathsf{R}})$ and $Dep_{\mathbf{pEff}}(\mathsf{A},{\mathsf{R}}^{\prime})$ is the same, i.e. the following diagram commutes
[…]
We conclude this section providing a direct correspondence between extensional dependent collections and internal groupoid action (see [MM12, JT94] for a precise definition) in $\mathbf{pEff}$. Indeed, for every object $(\mathsf{A},[\mathsf{R}])\in\mathbf{pEff}$, the kernel pair of a quotient arrow
$p:=[id_{\mathsf{A}}]:(\mathsf{A},[=_{\mathsf{A}}])\to(\mathsf{A},[\mathsf{R}])$
induces an internal groupoid $\mathsf{Ker}(p,{\mathsf{R}})$ obtained through the kernel pair of $p$
[…]
and for any other representative ${\mathsf{R}}^{\prime}\in[\mathsf{R}]$ the two internal groupoids are isomorphic in the category of internal groupoids of $\mathbf{pEff}$.
[…]
Proposition 6.11.
Let $(\mathsf{A},[\mathsf{R}])$ be an object in $\mathbf{pEff}$. There exists a full and faithful essentially surjective functor
[…]
Proof.
By general descent theory, there exists a canonical comparison functor (see [JT94, §2.2])
[…]
Since $p$ is a regular epimorphism and $\mathbf{pEff}$ is exact, then we obtain that $p$ is an effective descent morphism, i.e. the functor $\Phi^{p,{\mathsf{R}}}$ is full, faithful and essentially surjective (see [JT94, §2.4 ]). The functor $A$ is defined as the composition of $\mathbf{K}$ and $\Phi^{p,\mathsf{R}}$. ∎
7 The structure of fibred extensional sets on $\mathbf{pEff}$
Now we are going to investigate the categorical structure of extensional dependent sets in definition 6.6 over $\mathbf{pEff}$ as a subfibration of the codomain fibration of $\mathbf{pEff}$ defined as the Grothendieck construction of the pseudo-functor
$\mathbf{pEff}_{set}:\mathbf{pEff}^{\operatorname{op}}\to\mathbf{Cat}$
We are forced to deal with fibrations since families of collections and sets loose their functoriality under substitution when lifted over $\mathbf{pEff}$ to become “extensional". At this point, given that slice categories provide a natural account for the notion of families of collections over an object of $\mathbf{pEff}$ we will just work with subfibrations of the codomain fibration over $\mathbf{pEff}$.
For this purpose, we exploit the functor $\mathbf{K}$ of Proposition 6.8. In more detail, the action of $\mathbf{pEff}_{set}$ on an object $(\mathsf{A},[\mathsf{R}])$ is given by the full subcategory of $\mathbf{pEff}/(\mathsf{A},[\mathsf{R}])$ whose objects are arrows isomorphic to arrows in $\mathbf{K}(Dep_{\mathbf{pEff}_{set}}(\mathsf{A},{\mathsf{R}}))$ for (any) representative of $[\mathsf{R}]$, see Remark 6.10. If $[p]:(\mathsf{A}^{\prime},[\mathsf{R}^{\prime}])\to(\mathsf{A},[\mathsf{R}])$ is an arrow in $\mathbf{pEff}$, then the value of $\mathbf{Set}^{r}_{[p]}([f])$, for $[f]\in\mathbf{pEff}_{set}((\mathsf{A},[\mathsf{R}]))$ is given by the equivalence class of the pullback of $[f]$ along $[p]$ that will be denoted by $[p]^{*}[f]$.
To prove that the above pseudo-functor is well-defined we need the following lemma.
Lemma 7.1.
Let $[p]:(\mathsf{A}^{\prime},[\mathsf{R}^{\prime}])\to(\mathsf{A},[\mathsf{R}]),$ be an arrow in $\mathbf{pEff}$. If $[f]\in\mathbf{pEff}/(\mathsf{A},[\mathsf{R}])$ is isomorphic to the image trough $\mathbf{K}$ of an extensional dependent set $(\mathsf{B},[\mathsf{S}],\sigma)\in Dep_{\mathbf{pEff}_{set}}(\mathsf{A},{\mathsf{R}})$, then $[p]^{*}[f]$ is isomorphic to the image through $\mathbf{K}$ of an element in $Dep_{\mathbf{pEff}_{set}}(\mathsf{A},{\mathsf{R}^{\prime}})$.
Proof.
We assume that $[f]$ is isomorphic to $\mathbf{K}((\mathsf{B},[\mathsf{S}],\sigma))$, then we can choose two representative $f:\mathsf{A}^{\prime}\to\mathsf{A}$ in $\bm{\mathcal{C}}_{r}$ and $[g]:\mathsf{R}^{\prime}\to\mathsf{R}$ in $\overline{{\mathbf{Prop}^{r}_{\mathbf{s}}}}(\mathsf{A}\times\mathsf{A})$ (recall that $\overline{{\mathbf{Prop}^{r}_{\mathbf{s}}}}(\mathsf{A}\times\mathsf{A})$ is the poset reflection of $\mathbf{Set}^{r}(\mathsf{A}\times\mathsf{A})$) for $[f]$ and consider the extensional dependent set
$(\mathbf{Set}^{r}_{p}(\mathsf{B}),(\overline{{\mathbf{Prop}^{r}_{\mathbf{s}}}})_{\Sigma(p,\mathsf{B}\times\mathsf{B})}([\mathsf{S}]),\mathbf{Set}^{r}_{p\times p}(\sigma)\circ(\mathsf{id}\times g))$
in $Dep_{\mathbf{pEff}_{set}}(\mathsf{A},{\mathsf{R}^{\prime}})$ and observe that its image through $\mathbf{K}$ is isomorphic to the pullback $[p]^{*}[f]$. ∎
Remark 7.2.
Similar to Lemma 3.8, any arrow $[\mathsf{f}]$ in $\mathbf{pEff}_{set}(\mathsf{A},[\mathsf{R}])$ between two arrows isomorphic to the image through $\mathbf{K}$ of two objects in $(\mathsf{B},\mathsf{S},\sigma)$ and $(\mathsf{C},\mathsf{H},\eta)$ in $Dep_{\mathbf{pEff}_{set}}(\mathsf{A},\mathsf{R})$
$[\mathsf{f}]:\mathrm{dom}(\mathbf{K}(\mathsf{B},\mathsf{S},\sigma)\to\mathrm{dom}(\mathbf{K}(\mathsf{C},\mathsf{H},\eta))$
is itself in $\mathbf{pEff}_{set}(\mathrm{dom}(\mathbf{K}(\mathsf{C},\mathsf{H},\eta)))$.
Denoting with $\mathbf{pEff}_{set}^{\to}$ the Grothendieck construction of $\mathbf{pEff}_{set}$ we obtain the following commutative diagram of fibrations
${{{\mathbf{pEff}^{\to}_{set}}}}$${{{\mathbf{pEff}}^{\to}}}$${{\mathbf{pEff}}}$$\scriptstyle{\mathsf{cod}_{set}}$$\scriptstyle{\mathsf{cod}}$
We now study the categorical property of these fibrations. To do that, instead of working with slice categories over $(\mathsf{A},[\mathsf{R}])\in\mathbf{pEff}$ we exploit the equivalent categories of extensional dependent collections and sets and prove their properties. The following results do not depend on the chosen representative of $[\mathsf{R}]$ thanks to Lemma 6.7 and Remark 6.10.
Before moving to the main theorem of the section we provide a characterization of monomorphisms for families of collections, which will be instrumental to prove part of the theorem. It exploits a characterization that holds for suitable doctrines in general. Indeed, there exists a forgetful functor
$U:Dep_{\mathbf{pEff}}(\mathsf{A},\mathsf{R})\to\mathbf{pEff}$
which “forgets" the $\sigma$-component in a faithful way. It is obtained composing the functor $\mathbf{K}$ with the domain functor $\mathbf{pEff}/(\mathsf{A},[\mathsf{R}])\to\mathbf{pEff}$.
Lemma 7.3.
Let $(\mathsf{A},[\mathsf{R}])$ be an object of pEff. An arrow $[\mathsf{m}]:(\mathsf{B},[\mathsf{S}],\sigma)\rightarrow(\mathsf{C},[\mathsf{H}],\nu)$ in $Dep_{\mathbf{pEff}}(\mathsf{A},{\mathsf{R}})$ is a monomorphism if and only if
$[\mathsf{S}]=\overline{\mathbf{Prop}^{r}}_{\mathbf{I}(\mathsf{m}\times\mathsf{m})}([\mathsf{H}])$
Proof.
Since $U$ is faithful, it reflects monomorphisms. Hence, using [MR13b, Corollary 4.8] for the elementary quotient completion of ${(\overline{{\mathbf{Prop}^{r}}})}$ we obtain that the arrow
$U([\mathsf{m}])=[\mathbf{I}(\mathsf{m})]:(\Sigma(\mathsf{A},\mathsf{B}),\exists_{d_{\mathsf{R}}}\overline{{\mathbf{Prop}^{r}}}_{t_{\sigma}}[\mathsf{S}])\to(\Sigma(\mathsf{A},\mathsf{C}),\exists_{d_{\mathsf{R}}}\overline{{\mathbf{Prop}^{r}}}_{t_{\nu}}[\mathsf{H}])$
is a monomorphism if and only if
$\exists_{d_{\mathsf{R}}}\overline{{\mathbf{Prop}^{r}}}_{t_{\sigma}}[\mathsf{S}]=\overline{{\mathbf{Prop}^{r}}}_{\mathbf{I}(\mathsf{m})\times\mathbf{I}(\mathsf{m})}(\exists_{d_{\mathsf{R}}}\overline{{\mathbf{Prop}^{r}}}_{t_{\nu}}[\mathsf{H}])$
which is equivalent to
$[\mathsf{S}]=\overline{\mathbf{Prop}^{r}}_{\mathbf{I}(\mathsf{m}\times\mathsf{m})}([\mathsf{H}])$
by definition of the arrows $d_{\mathsf{R}}$ and $t_{\sigma}$ in the proof of Proposition 6.8.
Remark 7.4.
The same result can be obtained in this alternative way. Consider to lift the doctrine ${\mathbf{Prop}^{r}}$ over the slice of $\bm{\mathcal{C}}_{r}$ over $\mathsf{A}\in\bm{\mathcal{C}}_{r}$ as done in [Cio23]:
$(\overline{{\mathbf{Prop}^{r}}})_{/\mathsf{A}}:(\bm{\mathcal{C}}_{r}/\mathsf{A})^{\operatorname{op}}\rightarrow\mathbf{Pos}$
the action on an object $\mathsf{f}:\mathsf{X}\to\mathsf{A}$ of $\bm{\mathcal{C}}_{r}$ is given by ${\mathbf{Prop}^{r}}(\mathsf{X})$, and the action on an arrow from $\mathsf{f}$ to $\mathsf{g}:\mathsf{Y}\to\mathsf{A}$ (which is an arrow $\mathsf{h}:\mathsf{X}\to\mathsf{Y}$ such that $\mathsf{g}\circ\mathsf{h}=\mathsf{f}$) is given by ${\mathbf{Prop}^{r}}_{\mathsf{f}}$. Then, consider the elementary quotient completion of ${(\overline{{\mathbf{Prop}^{r}}})_{/\mathsf{A}}}$ and observe that there exists a forgetful functor
$U:Dep_{\mathbf{pEff}}(\mathsf{A},\mathsf{R})\to\mathcal{Q}_{(\overline{{\mathbf{Prop}^{r}}})_{/A}}$
which “forgets" the $\sigma$-component in a faithful way: it sends an object $(\mathsf{B},\mathsf{S},\sigma)$ to the pair $(\mathsf{p}_{1}^{\Sigma},\mathsf{S})$ in $\mathcal{Q}_{(\overline{{\mathbf{Prop}^{r}}})_{/A}}$, where $\mathsf{p}_{1}^{\Sigma}:\Sigma(\mathsf{A},\mathsf{B})\to\mathsf{A}$. Since, as shown in [Cio23], the elementary quotient completion of ${(\overline{{\mathbf{Prop}^{r}}})_{/A}}$ inherits some properties from $\overline{{\mathbf{Prop}^{r}}}$, and in particular it has full strict comprehensions and comprehensive diagonals, [MR13b, Corollary 4.8] applied to it and faithfulness of $U$ imply the same monomorphism characterization of lemma 7.3.
Theorem 7.5.
Let $(\mathsf{A},[\mathsf{R}])$ be an object of pEff. The categories $Dep_{\mathbf{pEff}}(\mathsf{A},\mathsf{R})$ and $Dep_{\mathbf{pEff}_{set}}(\mathsf{A},\mathsf{R})$ are locally cartesian closed list-arithmetic pretoposes.
Proof.
Since, $\mathbf{pEff}$ is equivalent to an exact completion of a lex category (see [MM21]) and $\mathbf{K}:Dep_{\mathbf{pEff}}(\mathsf{A},{\mathsf{R}})\rightarrow\mathbf{pEff}/(\mathsf{A},[\mathsf{R}])$ preserves and reflects all limits, colimits and exponentials, in the case of families of collection Theorem 7.5 follows from the properties of the exact completion. Indeed, exactness and local cartesian closure are preserved under slicing, such as extensivity (see [CLW93]) and parametrized list objects (see [Mai05]). However, when restricting to $\mathbf{pEff}_{set}(\mathsf{A},[\mathsf{R}])$ we must check that the above properties restrict to a proper subcategory of $\mathbf{pEff}/(\mathsf{A},[\mathsf{R}])$. Hence, we provide directly a proof of the result for $Dep_{\mathbf{pEff}}(\mathsf{A},\mathsf{R})$ to be able to restrict to $Dep_{\mathbf{pEff}_{set}}(\mathsf{A},\mathsf{R})$ and hence to $\mathbf{pEff}_{set}(\mathsf{A},[\mathsf{R}])$.
Finite limits.
  1. 1.
    A terminal object in $Dep_{\mathbf{pEff}}(\mathsf{A},\mathsf{R})$ is given by (1,,!)(1$\top$,!)( 1 , ⊤ , ! ). This is a terminal object also in $Dep_{\mathbf{pEff}_{set}}(\mathsf{A},\mathsf{R})$.
  2. 2.
    A binary product of $(\mathsf{B},[\mathsf{S}],\sigma)$ and $(\mathsf{C},[\mathsf{H}],\nu)$ is given by
    $(\mathsf{B}\times\mathsf{C},\overline{\textbf{Prop}^{r}}_{\textbf{I}(p_{1}\times p_{1})}([\mathsf{S}])\wedge\overline{\textbf{Prop}^{r}}_{\textbf{I}(p_{2}\times p_{2})}([\mathsf{H}]),\langle\sigma\circ(p_{1}\times\mathsf{id}_{\mathsf{R}}),\nu\circ(p_{2}\times\mathsf{id}_{\mathsf{R}})\rangle)$
    with projections inherited by those of $\mathsf{B}\times\mathsf{C}$. Since the product of two sets is a set and the conjunction of small propositions is a small proposition, the above construction restricts to $Dep_{\mathbf{pEff}_{set}}(\mathsf{A},\mathsf{R})$.
  3. 3.
    An equalizer of two arrows $[\mathsf{f}],[\mathsf{g}]:(\mathsf{B},[\mathsf{S}],\sigma)\rightarrow(\mathsf{C},[\mathsf{H}],\nu)$ is given by
    $(\Sigma^{\mathsf{A}}(\mathsf{B},\mathbf{Col}^{r}_{\mathbf{I}(\langle{\mathsf{f}},\mathsf{g}\rangle)}\mathsf{H}),\overline{\textbf{Prop}^{r}}_{\textbf{I}(\mathsf{p}_{1}^{\mathsf{A},\Sigma}\times\mathsf{p}_{1}^{\mathsf{A},\Sigma})}([\mathsf{S}]),\eta)$
    where $\eta$ acts on a realizer $r\Vdash\mathsf{R}(a,a^{\prime})$, a term $b:\mathsf{B}(a)$ and a realizer $q:\mathsf{H}_{a}(\mathsf{f}(a),\mathsf{g}(a))$, as $\sigma_{a,a^{\prime}}(r,b)$ and $q^{\prime}:\mathsf{H}_{a^{\prime}}(\mathsf{f}(\sigma_{a,a^{\prime}}(r,b)),\mathsf{g}(\sigma_{a,a^{\prime}}(r,b)))$, where $q^{\prime}$ is obtained through the concatenation of the term $q_{\nu}:\mathsf{H}_{a^{\prime}}(\nu_{a,a^{\prime}}(r,\mathsf{f}(b)),\nu_{a,a^{\prime}}(r,\mathsf{g}(b)))$ (from condition 1 of Definition 6.2 for $\nu$) and the terms $u:\mathsf{H}_{a^{\prime}}(\mathsf{f}(\sigma_{a,a^{\prime}}(r,b)),\nu_{a,a^{\prime}}(r,\mathsf{f}(b)))$ and $v:\mathsf{H}_{a^{\prime}}(\nu_{a,a^{\prime}}(r,\mathsf{g}(b)),\mathsf{g}(\sigma_{a,a^{\prime}}(r,b)))$ (from condition 2 of Definition 6.3 for $\mathsf{f}$ and $\mathsf{g}$ and symmetry).
Stable finite coproducts.
An initial object is given by $(0,\top,\sigma)$ where $\sigma$ is the unique possible one. A coproduct for $(\mathsf{B},[\mathsf{S}],\sigma)$ and $(\mathsf{C},[\mathsf{H}],\nu)$ is given by
$(\mathsf{B}+\mathsf{C},[\mathsf{S}]{+}[\mathsf{H}],(\sigma{+}\nu)j)$
where
  1. 1.
    $[\mathsf{S}]{+}[\mathsf{H}]=\exists_{\mathbf{I}(i_{1}\times i_{1})}([\mathsf{S}])\vee\exists_{\mathbf{I}(i_{2}\times i_{2})}([\mathsf{H}])$,
  2. 2.
    $j$ is the canonical isomorphism from $(\mathbf{Col}^{r}_{p_{1}}(\mathsf{B})+\mathbf{Col}^{r}_{p_{1}}(\mathsf{C}))\times{\mathsf{R}}$ to $(\mathbf{Col}^{r}_{p_{1}}(\mathsf{B})\times{\mathsf{R}})+(\mathbf{Col}^{r}_{p_{1}}(\mathsf{B})\times{\mathsf{R}})$ which exists since $\mathbf{Col}^{r}(\mathsf{A}\times\mathsf{A})$ is extensive.
together with the injections $[i_{1}]$ and $[i_{2}]$. One can check that these injections are mono, and that coproducts are stable under pullbacks using pullbacks constructed through the finite limits of $Dep_{\mathbf{pEff}}(\mathsf{A},\mathsf{R})$ and $Dep_{\mathbf{pEff}_{set}}(\mathsf{A},\mathsf{R})$ described above. Initial objects are trivially stable, since they are stable in $\mathbf{Col}^{r}(\mathsf{A})$.
These constructions provide stable finite coproducts also in $Dep_{\mathbf{pEff}_{set}}(\mathsf{A},\mathsf{R})$ using the same arguments together with Theorem 3.16 and Lemma 3.8.
List objects. A list object for $(\mathsf{B},[\mathsf{S}],\sigma)$ is given by $(\mathsf{List}(\mathsf{B}),[\mathsf{S}_{list}],\sigma_{list})$ where
  1. 1.
[…]
    $\mathsf{S}_{list}=\overline{{\mathbf{Prop}^{r}}}_{\mathbf{I}(\ell\times\ell)}(\exists_{\mathbf{I}(\Delta_{\mathsf{N}})}(\top))\wedge\qquad\qquad\qquad\qquad\qquad\qquad\qquad$
    $\forall_{\mathbf{I}(p_{1})}(\overline{{\mathbf{Prop}^{r}}}_{\mathbf{I}((\langle\mathsf{comp}\circ(p_{1}\times\mathsf{id}_{\mathsf{N}}))\times(\langle\mathsf{comp}\circ(p_{2}\times\mathsf{id}_{\mathsf{N}})))}(\exists_{\mathbf{I}(i_{1}\times i_{1})}([\mathsf{S}])\vee\exists_{\mathbf{I}(i_{2}\times i_{2})}(\top)))$
    where $\ell$ is the length arrow from $\mathsf{List}(\mathsf{B})\rightarrow\mathsf{N}$ in $\mathbf{Col}^{r}(\mathsf{A})$ and $\mathsf{comp}:\mathsf{List}(\mathsf{B})\times\mathsf{N}\rightarrow\mathsf{B}+1$ is the component arrow in $\mathbf{Col}^{r}(\mathsf{A})$ which selects the $n$-component of a list.
  2. 2.
    $\sigma_{list}$ is the unique arrow in $\mathbf{Col}^{r}(\mathsf{A})$ making the following diagram commute
    ${{{\mathsf{R}}}}$${{\mathsf{List}(\mathbf{Col}^{r}_{p_{1}}(\mathsf{B}))\times\mathsf{R}}}$${{(\mathsf{List}(\mathbf{Col}^{r}_{p_{1}}(\mathsf{B}))\times\mathbf{Col}^{r}_{p_{1}}(\mathsf{B}))\times\mathsf{R}}}$${{\mathsf{List}(\mathbf{Col}^{r}_{p_{2}}(\mathsf{B}))}}$${{\mathsf{List}(\mathbf{Col}^{r}_{p_{2}}(\mathsf{B}))\times\mathbf{Col}^{r}_{p_{2}}(\mathsf{B})}}$$\scriptstyle{\langle\epsilon\circ!,\mathsf{id}_{{\mathsf{R}}}\rangle}$$\scriptstyle{\epsilon\circ!}$$\scriptstyle{\sigma_{list}}$$\scriptstyle{\mathsf{cons}}$$\scriptstyle{\mathsf{cons}\times\mathsf{id}_{{\mathsf{R}}}}$$\scriptstyle{\langle\sigma_{list}\circ(p_{1}\times\mathsf{id}_{\mathsf{R}}),\sigma\circ(p_{2}\times\mathsf{id}_{\mathsf{R}})\rangle}$
These constructions can be performed also in $Dep_{\mathbf{pEff}_{set}}(\mathsf{A},\mathsf{R})$ thanks to Theorem 3.16 and Lemma 3.8.
Exactness.
Recall that $\mathbf{pEff}$ is exact and that the slice categories of an exact category are also exact. Hence, exactness of $Dep_{\mathbf{pEff}}(\mathsf{A},\mathsf{R})$ follows from the fact that $\mathbf{K}:Dep_{\mathbf{pEff}}(\mathsf{A},\mathsf{R})\to\mathbf{pEff}/(\mathsf{A},[\mathsf{R}])$ is a fully and faithful essentially surjective functor. To prove that $Dep_{\mathbf{pEff}_{set}}(\mathsf{A},\mathsf{R})$ is also exact, observe that the coequalizer of an equivalence relation on $(\mathsf{B},[\mathsf{S}],\sigma)\in Dep_{\mathbf{pEff}_{set}}(\mathsf{A},\mathsf{R})$
$[\mathsf{c}]:(\mathsf{C},[\mathsf{H}],\nu)\to(\mathsf{B},[\mathsf{S}],\sigma)\times(\mathsf{B},[\mathsf{S}],\sigma)$
is given by
$[\mathsf{id}_{\mathsf{B}}]:(\mathsf{B},[\mathsf{S}],\sigma)\to(\mathsf{B},[\mathsf{S}^{\prime}],\sigma)$
where
$[\mathsf{S}^{\prime}]:=\exists_{\mathbf{I}(\langle p_{2},p_{3}\rangle)}(\overline{{\mathbf{Prop}^{r}}}_{\mathbf{I}(\mathsf{c_{1}}\times p_{1})}([\mathsf{S}])\wedge\overline{{\mathbf{Prop}^{r}}}_{\mathbf{I}(\mathsf{c_{2}}\times p_{2})}([\mathsf{S}]))$
supposed $[\mathsf{c}_{1}]=p_{1}\circ[\mathsf{c}]$ and $[\mathsf{c}_{2}]=p_{2}\circ[\mathsf{c}]$.
The image through $\mathbf{K}$ of $[\mathsf{id}_{\mathsf{B}}]$
$[\mathbf{I}(\mathsf{id}_{\mathsf{B}})]:(\Sigma(\mathsf{A},\mathsf{B}),\exists_{d_{\mathsf{R}}}\overline{{\mathbf{Prop}^{r}}}_{t_{\sigma}}([\mathsf{S}]))\to(\Sigma(\mathsf{A},\mathsf{B}),\exists_{d_{\mathsf{R}}}\overline{{\mathbf{Prop}^{r}}}_{t_{\sigma}}([\mathsf{S}^{\prime}]))$
is a coequalizer in $\mathbf{pEff}$ of $\mathbf{K}([\mathsf{c}])$, that is in $\mathbf{pEff}_{set}(\Sigma(\mathsf{A},\mathsf{B}),\exists_{d_{\mathsf{R}}}\overline{{\mathbf{Prop}^{r}}}_{t_{\sigma}}([\mathsf{S}^{\prime}]))$ by Remark 7.2. Stability and effectiveness follow from the pullback property in Lemma 7.1.
Local cartesian closure. For two objects $(\mathsf{B},[\mathsf{S}],\sigma)$ and $(\mathsf{C},[\mathsf{H}],\nu)$ in $Dep_{\mathbf{pEff}}(\mathsf{A},\mathsf{R})$ consider the triple $(\mathsf{C}^{\mathsf{B}},[\mathsf{H}]^{[\mathsf{S}]},{\nu}^{\sigma})$ defined as follows. The object $\mathsf{C}^{\mathsf{B}}:=\Sigma^{\mathsf{A}}(\mathsf{B}\Rightarrow\mathsf{C},\Pi_{p_{1}}({\mathbf{Col}^{r}}_{\mathbf{I}(p_{2})}(\mathsf{S})\to{\mathbf{Col}^{r}}_{\mathbf{I}(\langle\mathsf{ev}\circ(\mathsf{id}\times p_{1}),\mathsf{ev}\circ(\mathsf{id}\times p_{2})\rangle)}(\mathsf{H}))$ is obtained through the weak exponential $\mathsf{B}\Rightarrow\mathsf{C}$, $\mathsf{ev}$ in $\mathbf{Col}^{r}(\mathsf{A})$. Intuitively, we restrict to those arrows $f_{a}\in\mathsf{B}(a)\Rightarrow\mathsf{C}(a)$ whose evaluation preserves $\mathsf{S}$ and $\mathsf{H}$, i.e. such that we have $s\Vdash(\forall b,b^{\prime}\in\mathsf{B}(a))\ (\mathsf{S}_{a}(b,b^{\prime})\to\mathsf{H}_{a}(f_{a}(b),f_{a}(b^{\prime})))$.
The relation $[\mathsf{H}]^{[\mathsf{S}]}$ is given by $\forall_{\mathbf{I}({p_{1}})}({\overline{{\mathbf{Prop}^{r}}}}_{\mathbf{I}(\langle\mathsf{ev}\circ(p_{1}\times\mathsf{id}),\mathsf{ev}\circ(p_{2}\times\mathsf{id})\rangle)}([\mathsf{H}]))$. Intuitively, we relate two arrows $f_{a},g_{a}\in\mathsf{B}(a)\Rightarrow\mathsf{C}(a)$ when we have $r\Vdash(\forall b\in\mathsf{B}(a))\ \mathsf{H}_{a}(f_{a}(b),g_{a}(b))$.
The action $\sigma^{\nu}$ is defined on a triple $((f_{a_{1}},s),r)$, where $r\Vdash\mathsf{R}(a_{1},a_{2})$, as the pair $(\tilde{f}_{a_{2}},\tilde{s})$ with $\tilde{f}_{a_{2}}(b):=\nu(f_{a_{1}}(\sigma(b,\mathsf{sym}(r))),r)$ and $\tilde{s}$ is obtained through $s$ and compatibility of $\sigma$ and $\nu$ with the corresponding equalities in condition (1) of Definition 6.2, and a witness $\mathsf{sym}$ of the symmetry of $\mathsf{R}$.
We now observe that the evaluation arrow $\mathsf{ev}:(\mathsf{B}\Rightarrow\mathsf{C})\times\mathsf{B}\to\mathsf{C}$ preserves $[\mathsf{H}]^{[\mathsf{S}]}$, $[\mathsf{H}]$ and $[\mathsf{S}]$. In the sense that, if $((f_{a},s),b)$ and $((f^{\prime}_{a},s^{\prime}),b^{\prime})$ are such that $r\Vdash\mathsf{H}_{a}(f_{a}(b),f^{\prime}_{a}(b))$ and $t\Vdash\mathsf{S}(b,b^{\prime})$, then using $s$ (applied to $t$), $r$ and the transitivity of $[\mathsf{H}]$ we obtain a $q\Vdash\mathsf{H}_{a}(f_{a}(b),f^{\prime}_{a}(b^{\prime}))$.
Hence, the evaluation arrow of $(\mathsf{C}^{\mathsf{B}},[\mathsf{H}]^{[\mathsf{S}]},{\nu}^{\sigma})$ is given by $[\mathsf{ev}]$ and we get an exponential object instead of a weak exponential as it happened in $\mathbf{Col}^{r}$.
Since weak exponentials of two sets are sets and small propositions are closed under connectives and quantification over sets, we obtain that the above construction restricts to $Dep_{\mathbf{pEff}_{set}}(\mathsf{A},\mathsf{R})$.
The local cartesian closure of $Dep_{\mathbf{pEff}}(\mathsf{A},\mathsf{R})$ follows observing that $\mathbf{K}:Dep_{\mathbf{pEff}}(\mathsf{A},\mathsf{R})\rightarrow\mathbf{pEff}/(\mathsf{A},[\mathsf{R}])$ induces a fully faithful and essentially surjective functor between the slice categories
$\mathbf{K}/(\mathsf{B},[\mathsf{S}],\sigma):Dep_{\mathbf{pEff}}(\mathsf{A},\mathsf{R})/(\mathsf{B},[\mathsf{S}],\sigma)\rightarrow(\mathbf{pEff}/(\mathsf{A},[\mathsf{R}]))/\mathbf{K}(\mathsf{B},[\mathsf{S}],\sigma)$
for every object $(\mathsf{B},[\mathsf{S}],\sigma)\in Dep_{\mathbf{pEff}}(\mathsf{A},\mathsf{R})$. Moreover, we have the isomorphism
$(\mathbf{pEff}/(\mathsf{A},[\mathsf{R}]))/\mathbf{K}(\mathsf{B},[\mathsf{S}],\sigma)\cong\mathbf{pEff}/\mathsf{dom}(\mathbf{K}(\mathsf{B},[\mathsf{S}],\sigma))$
and the fact that $\mathbf{pEff}$ is locally cartesian closed, implies that $Dep_{\mathbf{pEff}}(\mathsf{A},\mathsf{R})$ is locally cartesian closed by reflection of exponentials. To prove that also $Dep_{\mathbf{pEff}_{set}}(\mathsf{A},\mathsf{R})$ is locally cartesian closed, we observe that for every object $(\mathsf{B},[\mathsf{S}],\sigma)\in Dep_{\mathbf{pEff}_{set}}(\mathsf{A},\mathsf{R})$, $\mathbf{K}$ induces a fully faithful and essentially surjective functor
$\mathbf{K}/(\mathsf{B},[\mathsf{S}],\sigma):(Dep_{\mathbf{pEff}_{set}}(\mathsf{A},\mathsf{R}))/(\mathsf{B},[\mathsf{S}],\sigma)\rightarrow(\mathbf{pEff}_{set}(\mathsf{A},[\mathsf{R}]))/\mathbf{K}(\mathsf{B},[\mathsf{S}],\sigma)$
and we also have the isomorphism
$(\mathbf{pEff}_{set}(\mathsf{A},[\mathsf{R}]))/\mathbf{K}(\mathsf{B},[\mathsf{S}],\sigma)\cong\mathbf{pEff}_{set}(\mathsf{dom}(\mathbf{K}(\mathsf{B},[\mathsf{S}],\sigma)))$
by Remark 7.2, and $\mathbf{K}$ induces a fully faithful essentially surjective functor
$\mathbf{K}:Dep_{\mathbf{pEff}_{set}}(\mathsf{dom}(\mathbf{K}(\mathsf{B},[\mathsf{S}],\sigma))\to\mathbf{pEff}_{set}(\mathsf{dom}(\mathbf{K}(\mathsf{B},[\mathsf{S}],\sigma)))$
Hence, by preservation and reflection of exponentials we obtain that $Dep_{\mathbf{pEff}_{set}}(\mathsf{A},\mathsf{R})$ is locally cartesian closed.
8 The "small subobjects" classifier
Let $\mathbf{pEff}_{prop}:\mathbf{pEff}^{op}\to\mathbf{Pos}$ denote the elementary quotient completion of $\overline{{\mathbf{Prop}^{r}}}:\bm{\mathcal{C}}_{r}^{\operatorname{op}}\to\mathbf{Pos}$ whose value on an object $(\mathsf{A},[\mathsf{R}])$ is given by the descent data of $\overline{{\mathbf{Prop}^{r}}}(\mathsf{A})$, i.e.
$\mathbf{pEff}_{prop}(\mathsf{A},[\mathsf{R}]):=\left(\left\{[\mathsf{P}]\in\overline{{\mathbf{Prop}^{r}}}(\mathsf{A})\ |\ \;\overline{{\mathbf{Prop}^{r}}}_{p_{1}}([\mathsf{P}])\wedge[\mathsf{R}]\leq\overline{{\mathbf{Prop}^{r}}}_{p_{2}}([\mathsf{P}])\right\},\leq\right)$ (10)
where $p_{i}:\mathsf{A}\times\mathsf{A}\to\mathsf{A}$ are the projections for $i=1,2$, and whose action on arrows is given by that of $\overline{{\mathbf{Prop}^{r}}}$.
Let $\mathbf{pEff}_{props}:\mathbf{pEff}^{\operatorname{op}}\to\mathbf{Pos}$ denote the restriction of $\mathbf{pEff}_{prop}$ on the fibers of $\overline{{\mathbf{Prop}^{r}_{\mathbf{s}}}}:\bm{\mathcal{C}}_{r}^{\operatorname{op}}\to\mathbf{Pos}$. From [MM21] we recall the following result which appears as Theorem 5.7.
Theorem 8.1.
The functor $\mathbf{pEff}_{prop}$ is a first-order hyperdoctrine equivalent to the subobjects doctrine $\mathbf{pEff}_{prop}\cong\mathbf{Sub}_{\mathbf{pEff}}$. ∎
We now provide a similar correspondence for $\mathbf{pEff}_{props}$ regarding suitable subobjects. We will denote with $\mathbf{Sub}_{\mathbf{pEff}}^{\mathbf{s}}$ the doctrine which sends an object $(\mathsf{A},[\mathsf{R}])\in\mathbf{pEff}$ to the poset of subobjects of $\mathbf{pEff}$ over $(\mathsf{A},[\mathsf{R}])$ which have a representative in $\mathbf{pEff}_{set}^{\to}$. This is a well-defined functor thanks to Lemma 7.1. We will refer to these objects as small subobjects.
Lemma 8.2.
Let $(\mathsf{A},\mathsf{R})$ be an object of pEff. Given an element $(\mathsf{B},[\mathsf{S}],\sigma)\in Dep_{\mathbf{pEff}}(\mathsf{A},[\mathsf{R}])$, $\mathbf{K}(\mathsf{B},[\mathsf{S}],\sigma)$ is a monomorphism in $\mathbf{pEff}$ if and only if $[\mathsf{S}]=\top$.
Proof.
This follows from Lemma 7.3, and the fact that an arrow $m$ in $\mathbf{pEff}$ with codomain $(\mathsf{A},[\mathsf{R}])$ is a monomorphism if and only if it is a monomorphism in $\mathbf{pEff}/(\mathsf{A},[\mathsf{R}])$ seen as an arrow from $m$ to $id_{(\mathsf{A},[\mathsf{R}])}$. Indeed, $id_{(\mathsf{A},[\mathsf{R}])}$ is isomorphic to $\mathbf{K}(1,\top,!)$. ∎
Proposition 8.3.
The functor $\mathbf{pEff}_{props}$ is equivalent to the small subobjects doctrine $\mathbf{Sub}_{\mathbf{pEff}}^{\mathbf{s}}$.
Proof.
Given an object $(\mathsf{A},[\mathsf{R}])\in\mathbf{pEff}$, consider a subobject over it which has a representative of the form $\mathbf{K}(\mathsf{B},[\mathsf{S}],\sigma)$ where $(\mathsf{B},[\mathsf{S}],\sigma)\in Dep_{\mathbf{pEff}_{set}}(\mathsf{A},\mathsf{R})$. We can associate to it the element $[\mathsf{B}]\in\overline{{\mathbf{Prop}^{r}_{\mathbf{s}}}}(\mathsf{A})$. Thanks to $\sigma$, it follows that $[\mathsf{B}]\in\mathbf{pEff}_{props}(\mathsf{A},[\mathsf{R}])$. Moreover, for two monomorphisms of this form, it holds that $\mathbf{K}(\mathsf{B},[\mathsf{S}],\sigma)\leq\mathbf{K}(\mathsf{B}^{\prime},[\mathsf{S}^{\prime}],\sigma^{\prime})$ if and only if $[\mathsf{B}]\leq[\mathsf{B}^{\prime}]$ in $\mathbf{pEff}_{props}(\mathsf{A},[\mathsf{R}])$. Conversely, given a $[\mathsf{P}]\in\mathbf{pEff}_{props}(\mathsf{A},[\mathsf{R}])$ and a representative $\sigma:\mathbf{Col}^{r}_{p_{1}}(\mathsf{P})\times\mathsf{R}\to\mathbf{Col}^{r}_{p_{2}}(\mathsf{P})$ (whose existence is guaranteed by (10)), we associate the equivalence class of $\mathbf{K}(\mathsf{P},\top,\sigma)$ which is a monomorphism thanks to Lemma 8.2. This provides the desired equivalence. ∎
Applying the same argument in [MM21, Thm. 5.12], and combining it with the previous proposition, we obtain a classifier for small subobjects.
Theorem 8.4.
There exists an object $\Omega$ which represents $\mathbf{Sub}_{\mathbf{pEff}}^{\mathbf{s}}$, i.e. there exists a natural isomorphism between the functors $\mathbf{Sub}_{\mathbf{pEff}}^{\mathbf{s}}(-)$ and $\mathbf{pEff}(-,\Omega)$. ∎
We now summarize the main results obtained in this work.
Theorem 8.5.
The category $\mathbf{pEff}$ is equipped with two fibrations
${{{\mathbf{pEff}^{\to}_{set}}}}$${{{\mathbf{pEff}}^{\to}}}$${{\mathbf{pEff}}}$$\scriptstyle{\mathsf{cod}_{set}}$$\scriptstyle{\mathsf{cod}}$
satisfying the following properties:
  1. 1.
    $\mathbf{pEff}$ is a locally cartesian closed list arithmetic pretopos;
  2. 2.
    for each $(\mathsf{A},[\mathsf{R}])\in\mathbf{pEff}$ the fiber $\mathsf{cod}_{set}^{-1}(\mathsf{A},[\mathsf{R}])$ is a locally cartesian closed list arithmetic pretopos;
  3. 3.
    for every arrow $[f]:(\mathsf{A},[\mathsf{R}])\to(\mathsf{A}^{\prime},[\mathsf{R}^{\prime}])$ in $\mathbf{pEff}$, the functor
    $[f]^{*}:\mathsf{cod}_{set}^{-1}(\mathsf{A}^{\prime},[\mathsf{R}^{\prime}])\to\mathsf{cod}_{set}^{-1}(\mathsf{A},[\mathsf{R}])$
    preserves the locally cartesian closed list-arithmetics pretopos structure;
  4. 4.
    for each $(\mathsf{A},[\mathsf{R}])\in\mathbf{pEff}$ the inclusion $\mathsf{cod}_{set}^{-1}(\mathsf{A},[\mathsf{R}])\to\mathbf{pEff}/(\mathsf{A},[\mathsf{R}])$ preserves the locally cartesian closed list-arithmetics pretopos structure;
  5. 5.
    there exists an object $\Omega$ classifying $\mathbf{Sub}_{\mathbf{pEff}}^{\mathbf{s}}$;
  6. 6.
    for every $[f]\in\mathsf{cod}_{set}^{-1}(\mathsf{A},[\mathsf{R}])$ there exists an exponential object $(\pi_{\Omega})^{[f]}$ in $\mathbf{pEff}/(\mathsf{A},[\mathsf{R}])$ where $\pi_{\Omega}:(\mathsf{A},[\mathsf{R}])\times\Omega\to(\mathsf{A},[\mathsf{R}])$ is the first projection;
  7. 7.
    $\mathbf{pEff}$ satisfies the formal Church’s thesis (see [MM21, Theorem 5.8]).
Proof.
1. follows from Theorem 5.1, 2. follows from Theorem 8.1, 3.  4.  5. and 6. follow from the constructions defined to prove Theorem 7.5. 7. follows from Theorem 8.4. ∎
The above properties testify that $\mathbf{pEff}$ is a fibred predicative variant of Hyland’s Effective Topos. Indeed, the above properties are validated after substituting $\mathbf{pEff}$ with ${\mathbf{Eff}}$ and identified fibred sets with all the objects of a slice category of ${\mathbf{Eff}}$.
9 Towards a direct interpretation of $\mathbf{emTT}$ into $\mathbf{pEff}$
The realizability interpretation of the intensional level of the Minimalist Foundation in [IMMS18, MMR21, MMR22] can be turned into a morphism of doctrines from the syntactical doctrine $G^{\mbox{{\bf mTT}}}$ arising from mTT (as introduced in [MR13b] sec.7.2) to the doctrine $\overline{{\mathbf{Prop}^{r}}}$, as represented on the left side of the following diagram.
${{\mathcal{CM}^{\operatorname{op}}}}$${{\bm{\mathcal{C}}_{r}^{\operatorname{op}}}}$${{\mathcal{Q}_{G^{\mathbf{mTT}}}^{\operatorname{op}}}}$${{\mathbf{pEff}^{\operatorname{op}}}}$${{\mathbf{Pos}}}$${{\mathbf{Pos}}}$$\scriptstyle{G^{\mathbf{mTT}}}$$\scriptstyle{{\overline{\mathbf{Prop}^{r}}}\cong\mathbf{wSub}_{\bm{\mathcal{C}}_{r}}}$$\scriptstyle{\overline{(G^{\mathbf{mTT}})}}$$\scriptstyle{\mathbf{pEff}_{prop}\cong\mathbf{Sub}_{\mathbf{pEff}}}$
Such a morphism lifts to a morphism between the corresponding elementary quotient completions, as represented on the right side of the above diagram. In the category $\mathcal{Q}_{G^{\mathbf{mTT}}}$ one can define an adequate notion of canonical isomorphisms, as done in [Mai09], thus producing a model for $\mathbf{emTT}$. This, combined with the morphism above and the fibered structure of $\mathbf{pEff}$ provided in this article, makes possible a direct interpretation of $\mathbf{emTT}$, also extended with inductive and coinductive predicates [MMR21, MMR22, MS23, Sab24], within $\mathbf{pEff}$.
10 Comparison with related works
In [BM11], van den Berg and Moerdijk propose a construction of a predicative rendering of Hyland’s Effective Topos in the wider context of algebraic set theory. Starting from a category with small maps $(\mathcal{E},\mathcal{S})$ they first construct a category of assemblies $\mathcal{A}sm_{\mathcal{E}}$ endowed with a family of maps induced on it by $\mathcal{S}$, and identified its exact completion $(\mathcal{Eff}_{\mathcal{E}},\mathcal{S}_{\mathcal{E}})$ as their notion of predicative realizability category (where $\mathcal{Eff}_{\mathcal{E}}$ is a suitable subcategory of the exact on regular completion of $\mathcal{A}sm_{\mathcal{E}}$). We can compare our construction with theirs, as soon as we take for $\mathcal{E}$, the category $\mathcal{E}[\mathbf{T}]$ of definable classes in an extension $\mathbf{T}$ of $\mathbf{CZF}$ having the numerical existence property, endowed with the collection $\mathcal{S}[\mathbf{T}]$ of set-fibered functional relations between them (see sec.8 of [BM08]). Examples of such a theory $\mathbf{T}$ are $\mathbf{CZF}$ itself and $\mathbf{CZF}+\mathbf{REA}$ as proven in [Rat05] and also $\mathbf{CZF}+\mathbf{RDC}+\bigcup-\mathbf{REA}$ (by adequately modifying the proof of Theorem 7.2 in [Rat05] and combining it with Theorem 7.4 in [Rat08]).
Over such a category, assemblies are pairs $(A,\alpha)$ in which $A$ is a definable class of $\mathbf{T}$ and $\alpha$ is a definable relation between elements of $A$ and of $\mathbb{N}$ such that $\mathbf{T}\vdash\forall x\in A\,\exists y\in\mathbb{N}\,\alpha(x,y)$. Arrows between assemblies $(A,\alpha)$ and $(B,\beta)$ are equivalence classes of functional relations $F$ from $A$ to $B$ such that
$\mathbf{T}\vdash\exists e\in\mathbb{N}\,\forall x\in A\,\forall y\in B\,\forall z\in\mathbb{N}\,(\alpha(x,z)\wedge F(x,y)\rightarrow\beta(y,\{e\}(z)))$
The assemblies category is denoted in [BM11] as $\mathcal{Asm}_{\mathcal{E}[\mathbf{T}]}$.
If we now consider our category $\bm{\mathcal{C}}_{r}$ of realized collections in section. 3.1 formalized in $\mathbf{T}$, every object $\{x|\,\varphi(x)\}$ (for which $\mathbf{T}\vdash\forall x\,(\varphi(x)\rightarrow x\in\mathbb{N})$ always holds) can be mapped to the assembly $\mathbf{i}(\{x|\,\varphi(x)\}):=(\{x|\varphi(x)\},=_{\varphi})$ where
$x=_{\varphi}y\equiv^{def}\varphi(x)\wedge\varphi(y)\wedge x=y$
Arrows $[\mathbf{n}]_{\approx}$ in $\bm{\mathcal{C}}_{r}$ can be sent to the functional relations representing their graph, thus defining a functor $\mathbf{i}$ from $\bm{\mathcal{C}}_{r}$ to $\mathcal{Asm}_{\mathcal{E}[\mathbf{T}]}$. Since the numerical existence property holds for $\mathbf{T}$, the functor $\mathbf{i}$ is full and faithful, thus $\bm{\mathcal{C}}_{r}$ can be identified as a full subcategory of $\mathcal{Asm}_{\mathcal{E}[\mathbf{T}]}$. Now, let us call recursive objects those assemblies where $A$ is a subclass of natural numbers satisfying
$\mathbf{T}\vdash\forall x\in A\,\forall y\in\mathbb{N}\,(\alpha(x,y)\leftrightarrow x=y)$
[…]
${{\mathbf{pEff}^{\to}_{set}}}$${{\mathbf{pEff}^{\to}}}$${{\mathcal{Disc}_{\mathcal{E}[\mathbf{T}]}^{\to}}}$${\mathbf{pEff}}$${{\mathcal{Disc}_{\mathcal{E}[\mathbf{T}]}}}$
In addition, recalling that $\mathcal{Disc}_{\mathcal{E}[\mathbf{T}]}$ is the ex/reg completion of the category of modest sets (see [Ros90]), namely the category of assemblies which satisfy
$\mathbf{T}\vdash\forall x,y\in A\forall z\in\mathbb{N}(\alpha(x,z)\wedge\alpha(y,z)\rightarrow x=y)$
[…]
Proposition 10.1.
Let $\mathbf{pEff}$ be the predicative effective topos built on an extension $\mathbf{T}$ of $\mathbf{CZF}$ having the numerical existence property; and let $\mathbf{pEff}_{sm.\Delta}$ be the full subcategory of $\mathbf{pEff}$ given by objects $(\mathsf{A},[\mathsf{R}])$ where $[\mathsf{R}]\in{\mathbf{Prop}^{r}_{\mathbf{s}}}(\mathsf{A}\times\mathsf{A})$. Then, $\mathbf{pEff}_{sm.\Delta}$ is a full subcategory of the predicative realizability category $\mathcal{Eff}_{\mathcal{E}[\mathbf{T}]}$ defined as in [BM11] from the syntactic category $\mathcal{E}[\mathbf{T}]$ of definable classes of $\mathbf{T}$ endowed with the family $\mathcal{S}[\mathbf{T}]$ of arrows whose fibers are sets in $\mathbf{T}$.
[…]
Proof.
The category $\mathcal{Eff}_{\mathcal{E}[\mathbf{T}]}$ proposed in [BM11] as a predicative version of Hyland’s Effective Topos ${\mathbf{Eff}}$ is the full subcategory of the ex/reg completion of $\mathcal{Asm}_{\mathcal{E}[\mathbf{T}]}$, whose objects are separated in the sense of pg. 141 of [BM08]. The statement follows observing that objects $(\mathsf{A},[\mathsf{R}])$ of $\mathbf{pEff}_{sm.\Delta}$, seen as objects in $(\mathcal{Asm}_{\mathcal{E[\mathbf{T}]}})_{ex/reg}$, are separated in the above sense since $[\mathsf{R}]\in{\mathbf{Prop}^{r}_{\mathbf{s}}}(\mathsf{A}\times\mathsf{A})$. ∎
Observe that, when we consider $\mathbf{T}$ to be $\mathbf{ZFC}$, although it does not satisfy the numerical existence property, we can still define a functor from $\mathbf{pEff}$ to a version ${\mathbf{Eff}}_{def}$ of Hyland’s Effective Topos built as the ex/reg completion of the full subcategory of $\mathcal{Asm}_{\mathcal{E}[\mathbf{ZFC}]}$ whose objects are assemblies of which the underlying classes are definable sets. Indeed, in $\mathbf{ZFC}$ all definable subclasses of $\mathbb{N}$ are in fact sets. In the same framework, one can see that this ${\mathbf{Eff}}_{def}$ is a full subcategory of $\bf\mathcal{Eff}_{\mathcal{E}[\mathbf{ZFC}]}$.
${\mathbf{pEff}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{not\ full}$${{\mathbf{Eff}}_{def}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{Eff}_{\mathcal{E}[\mathbf{ZFC}]}}$
A comparison between $\mathbf{pEff}$ formalized in $\widehat{ID_{1}}$, and the usual Hyland’s Effective Topos ${\mathbf{Eff}}$ as an internal large category within ZFC is induced by the standard interpretation of $\widehat{ID_{1}}$ in the ZFC set theory, as shown in [MM21].
All these observations have taken into account only the base category $\mathbf{pEff}$. A comparison between their fibered structures is left to future work. Just observe that while maps in van den Berg and Moerdijk ’s approach are built on the notion of set in axiomatic set theory, in our approach the fibration of sets is defined mimicking the structure of sets within the language of dependent type theory as that used in both levels of the Minimalist Foundation [Mai09].
Nevertheless, our comparisons are sufficient to conclude that the full subcategory of discrete objects of Hyland’s Effective Topos or of van den Berg and Moerdijk’s constructive rendering contains already a predicative version of $\mathbf{Eff}$. In a few words, here we confirm what said on p. 1140 of [vO15] by concluding that the discrete objects of $\mathbf{Eff}$ can be considered the realm of constructive predicative mathematics.
11 Conclusions
We have shown here that the full subcategory of discrete objects of Hyland’s Effective topos $\mathbf{Eff}$ contains already a fibred predicative topos validating the formal Church’s thesis, even when both are formalized in a constructive metatheory. We did so by enucleating a fibred structure for $\mathbf{pEff}$ introduced in [MM21] that would be enough to model both levels of MF and its extensions with inductive and conductive predicates in [MMR21, MMR22, MS23, Sab24].
Furthermore, thanks to the equiconsistency of both levels of MF with their classical version shown in [MS24], we could use $\mathbf{pEff}$ as a computational model for the classical version of MF, too. And, since both levels of MF can be interpreted in $\mathbf{HoTT}$ in a compatible way as shown in [CM24], it would be interesting to investigate whether $\mathbf{pEff}$ is somewhat related to computational models of Homotopy Type theory, for example to that in [SA21].
In the future, we intend to investigate abstract notions of fibred predicative topos and of predicative tripos-to-topos construction of which the fibred structure of $\mathbf{pEff}$ enucleated here would provide a key example.
Acknowledgements The authors would like to acknowledge Michael Rathjen, Pino Rosolini, Pietro Sabelli, Thomas Streicher and Davide Trotta for fruitful discussions.
All the authors are affiliated with the INdAM national research group GNSAGA. Moreover, the first and second authors would like to thank the Hausdorff Research Institute for Mathematics for the financial support received to attend part of the trimester program ’Prospects of Formal Mathematics’ when significant progress was made for this work.